### v1

# Friend or Foe? Harnessing Controllable Overfitting for Anomaly Detection

Abstract
Overfitting has long been stigmatized as detrimental to model performance, especially in the context of anomaly detection. Our work challenges this conventional view by introducing a paradigm shift—recasting overfitting as a controllable and strategic mechanism for enhancing model discrimination capabilities. In this paper, we present Controllable Overfitting-based Anomaly Detection (COAD), a novel framework designed to leverage overfitting for optimized anomaly detection. We propose the Aberrance Retention Quotient (ARQ), a novel metric that systematically quantifies the extent of overfitting, enabling the identification of an optimal ”golden overfitting interval” $ARQ_{optimal}$. Within this interval, overfitting is leveraged to significantly amplify the model’s sensitivity to anomalous patterns, while preserving generalization to normal samples. Additionally, we present the Relative Anomaly Distribution Index (RADI), an innovative metric designed to complement AUROC-pixel by providing a more versatile and theoretically robust framework for assessing model performance. RADI leverages ARQ to track and evaluate how overfitting impacts anomaly detection, offering an integrated approach to understanding the relationship between overfitting dynamics and model efficacy. Our theoretical work also rigorously validates the use of Gaussian noise in pseudo-anomaly synthesis, providing the foundation for its broader applicability across diverse domains. Empirical evaluations demonstrate that our controllable overfitting method not only achieves State-Of-The-Art(SOTA) performance in both one-class and multi-class anomaly detection tasks but also redefines overfitting from a modeling challenge into a powerful tool for optimizing anomaly detection.
1 Introduction
In the traditional context of machine learning, overfitting has long been perceived as an undesirable phenomenon, just like  [3, 8, 10, 20], traditionally viewed as the result of a model excessively memorizing training data, leading to compromised generalization capabilities. However, we challenge this conventional wisdom by extending the concept beyond the domain of anomaly detection, positioning controllable overfitting as a versatile mechanism for enhancing model sensitivity across a wide range of applications. This leads us to the central question that underpins our work: can overfitting be a friend rather than a foe in the domain of anomaly detection?
(a) False Positive Cases
(b) False Negative Cases
Figure 1: Each subfigure presents a sequence from left to right, displaying: the original image, the ground truth, the results from other methods, like RD [6], RD++[23], UniAD [28] and DiAD [11], and our results after overfitting enhancement based on existing frameworks with COAD. (a) illustrates a significant reduction in the false positive rate after applying our overfitting approach. (b) highlights the reduction in the false negative rate, demonstrating enhanced sensitivity towards true anomalies with our method. The consistent improvement across both metrics indicates the effectiveness of our methods in refining detection accuracy.
The conventional understanding of overfitting necessitates a significant paradigm shift. Historically stigmatized as inherently detrimental to machine learning models, overfitting, as our research demonstrates, can under specific controlled conditions be repurposed as a powerful mechanism for enhancing model performance. In scenarios where training data consists solely of normal samples, we show that allowing the model to comprehensively memorize these features enhances its sensitivity to deviations—namely, anomalies—during testing. This reframing positions overfitting not as an inherent flaw to be avoided, but as a strategic asset to be leveraged effectively.
Anomaly detection in visual domains is a crucial task that aims to identify patterns that do not conform to expected behavior. Traditionally, anomaly detection relies on the notion of learning the distribution of normal data and distinguishing anomalies by identifying outliers. Various methods have been proposed to achieve this, including reconstruction-based approaches [31, 15, 30], embedding-based models [25, 32, 13, 9], and anomaly synthesis technique [4, 28, 17, 23].
We advance the notion of controllable overfitting, suggesting that a calculated and moderate degree of overfitting enables the model to capture intricate and subtle characteristics of the data. This heightened focus allows for a pronounced contrast when confronted with previously unseen deviations—precisely what is required for successful detection tasks, as opposed to a simplistic pursuit of generalization. This approach provides an opportunity to redefine how we evaluate model capabilities in environments devoid of training data that represents the full spectrum of possible scenarios. In this paper, we present Controllable Overfitting-based Anomaly Detection, called COAD.
Our theoretical framework and empirical validation jointly support this proposition. We demonstrate that, as the degree of overfitting increases, the score distribution for normal samples evolves accordingly, while the score distribution for anomalous samples remains stable. By introducing the Aberrance Retention Quotient, called ARQ, as an overfitting parameter, coupled with a noise loss term to finely control the degree of overfitting, we have identified an optimal ”Golden Overfitting Interval” $ARQ_{optimal}$ that maximizes detection efficacy without compromising model reliability. This strategic balance enables robust differentiation between normal and anomalous samples, providing a substantial performance boost.
We also introduce a novel metric, the Relative Anomaly Distribution Index, called RADI, to differentiate from traditional AUROC-pixel measures. Unlike AUROC-pixel, RADI is derived from the cumulative distribution function (CDF) of the normal and anomalous pixel prediction score distributions. While RADI provides an indirect reflection of detection performance and maintains a similar trend to AUROC-pixel, it also offers a more versatile framework for theoretical modeling and practical discussion of model capabilities. This distinction underscores its value as a complementary tool, adding a new layer of analytical depth to our approach.
Moreover, we substantiate the rationale for employing Gaussian noise in pseudo-anomaly detection modules. Through calculating the Total Variation Distance (TVD) between actual anomaly data and a Gaussian distribution across the entire dataset, we obtain a fair result. This outcome indicates a high level of similarity between true anomalies and a Gaussian distribution, thereby providing a theoretical validation for the common practice of using Gaussian noise as an effective pseudo-anomaly generator in previous works like GLASS [4] and SimpleNet [17]. This insight not only strengthens our methodology but also offers a theoretical foundation for broader applications in anomaly detection frameworks.
To summarize, the key contributions of our work are as follows:
  • We challenge and transcend the conventional understanding of overfitting, positioning it as a controllable and transformative mechanism capable of unlocking model capabilities beyond traditional boundaries.
  • We introduce the ARQ to precisely regulate the overfitting degree and the RADI as a complementary metric to AUROC-pixel, providing a more adaptable tool for mathematical modeling and facilitating Dual Control Mechanism with the use of COAD.
  • Our approach dismantles the demonization of overfitting, repurposing it as a generalizable module and achieving SOTA results in both one-class and multi-class anomaly detection tasks. We also provide a theoretical foundation for utilizing Gaussian noise as a preliminarily pseudo-anomaly generator.
2 Related Work
2.1 Overview of Anomaly Detection
Reconstruction-Based Method
Reconstruction-based methods are founded on the hypothesis that models trained exclusively on normal data can successfully reconstruct normal instances while failing to do so for anomalies, thereby using reconstruction error as a measure of abnormality. Notable reconstruction-based approaches include Autoencoders (AE)[29, 33], Generative Adversarial Networks (GANs)[7, 15], and Reverse Distillation (RD)[6, 23]. Specific methods under these categories, such as DSR[31], RealNet[32], and DRAEM[30], attempt to learn the representation of normal data and generate accurate reconstructions for normal samples but tend to produce significant residuals for anomalous inputs.
Embedding-Based Method
Recent developments have seen the use of embedding-based methods, besides reconstruction-based methods. These models leverage pre-trained networks to extract features from input images, aiming to separate anomalies from normal samples within the feature space. Anomalies are identified based on the distance between their embeddings and those of normal samples. Techniques such as Double-MMD RGP[25], RealNet[32], PyramidFlow[13], and CFlowAD[9] utilize this paradigm, some of which employ a memory bank to store representative features and apply distance metrics to detect anomalies. Embedding-based methods are particularly valuable in high-dimensional feature spaces, where the intricate relationships between different regions of the data can be better captured using a pre-trained model’s learned representations.
Synthetic Anomaly Generation
Another line of research focuses on synthetic anomaly generation, a strategy that involves creating pseudo-anomalies to augment training data and improve detection performance. These synthetic anomalies are used to provide explicit anomaly labels in the training set, converting the unsupervised task into a form of supervised learning. Notable techniques include CutPaste [14], which pastes cut-out normal patches in different positions, and methods such as GLASS[4], UniAD[28], SimpleNet[17], RealNet[32],and DDPM [12], all of which employ Gaussian noise to generate synthetic anomalies at both feature and image levels. These approaches help in training the model to differentiate between normal and abnormal samples, thus improving its ability to detect true anomalies. By contrast, methods such as DRAEM[30] and DMDD [16] employ Berlin noise or other forms of noise, which have not demonstrated the same level of effectiveness in anomaly detection tasks. In subsequent sections, we will theoretically justify the use of Gaussian noise in these pseudo-anomaly generation modules, establishing its effectiveness and reasonableness as an anomaly synthesis strategy.
2.2 One-class vs. Multi-class
Our method builds on existing frameworks in anomaly detection and requires a deep understanding of these foundational models. Without a comprehensive grasp of their structure and mechanisms, our method may be misapplied, leading to suboptimal outcomes. Hence, we will conduct demonstrative experiments from both one-class and multi-class perspectives.
One-class Anomaly Detection
One-class anomaly detection focuses on modeling the distribution of normal data to identify deviations. Methods such as RD [6] and RD++ [23] are foundational in this area. RD employs a pretrain model for feature extraction, while RD++ further enhances this process for improved robustness.
Multi-class Anomaly Detection
Multi-class anomaly detection handles diverse object categories and intra-class variations. Notable approaches like UniAD [28] and HVQ-Trans [18] effectively generalize across multiple classes.
2.3 Non-Diffusion vs. Diffusion-Based
For the purpose of making our method easier for readers to understand and apply, we have also supplemented it with the diffusion-based framework.
Non-diffusion methods
Non-diffusion methods, such as UniAD [28], have been effective for anomaly detection but often face limitations in preserving the semantic consistency of anomalous regions, especially in complex settings.
Diffusion-based methods
Diffusion-based methods have recently gained traction due to their exceptional reconstruction capabilities, as evidenced by DiAD [11], which stands out among contemporary methods. DiAD employs a diffusion framework for generating anomalies while maintaining semantic information.
3 Method
3.1 Overview
(a) Training Stage
(b) Inference Stage
Figure 2: Comprehensive Framework for COAD.(a) Training Stage: In the traditional framework, the teacher model remains frozen throughout. The pseudo-anomaly generation and filtering components are optional, remaining unfrozen during the normal training stage but frozen during the controllable overfitting stage, under the guidance of the Dual Control Mechanism. The student model also remains unfrozen during normal training stage, with selective freezing employed during the controllable overfitting stages. When the Dual Control Mechanism fails, Freeze Command Signals are triggered, prompting the selective freezing of certain network layers in the traditional framework to effectively manage overfitting. (b) Inference Stage: The inference pipeline follows a similar procedure to the traditional framework, ensuring that the enhancements provided by COAD integrate seamlessly.
As  Fig. 2 shows, our works provides an overview of our proposed framework for enhancing anomaly detection through controllable overfitting. We introduce Controllable Overfitting-based Anomaly Detection, called COAD. We systematically integrate novel elements such as ARQ for quantifying overfitting and RADI as a complementary metric for AUROC-pixel. These components collectively facilitate the monitoring, control, and optimization of the overfitting process to maximize model sensitivity while preserving generalization capabilities.
3.2 Quantifying Overfitting with Aberrance Retention Quotient
The exploration of overfitting has largely been confined to its avoidance, and there has been a conspicuous lack of systematic studies on how to control or even beneficially leverage overfitting for enhancing model capabilities. Moreover, existing research lacks a quantitative metric that accurately reflects the extent of overfitting, which is critical for understanding its implications and finding the ”sweet spot” in model training. To fill this gap, we introduce the Aberrance Retention Quotient , called ARQ, a novel metric designed to quantify the degree of overfitting by capturing the model’s divergence from true data representation during training.
The ARQ is formally defined as follows:
$ARQ=\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})}{\sum_{i=1}^{N}y_{i}},$ (1)
where N represents the total number of instances, $\hat{y}_{i}$ represents the predicted output at the $i^{th}$ instance, and $y_{i}$ denotes the corresponding original ground truth value. The numerator captures the aggregate prediction deviation from the true labels across all data points, while the denominator represents the totality of the predicted target values, normalizing the aberrance in the context of the entire dataset.
ARQ is a core metric in our framework that quantifies the progression of overfitting during model training. By tracking ARQ, we identify the golden overfitting interval, where the model’s sensitivity to anomalous patterns is maximized without compromising generalization. Specifically, we denote the interval of optimal ARQ as:
$ARQ_{\text{optimal}}=[\theta-\delta,\theta+\delta],$ (2)
where $\theta$ is the baseline value around which the optimal range is established, $\delta$ is a quantity of the same order of magnitude as $\theta$, representing the permissible deviation from the baseline $\theta$ within the optimal range, which indicates the region in which the overfitting is leveraged most effectively.
As overfitting progresses, $ARQ_{optimal}$ helps us mitigate false positives while reducing false negatives,  Fig. 1(a) and  Fig. 1(b) illustrate some cases. By effectively using ARQ, we transform overfitting into a controlled mechanism that enhances model robustness and discriminative power for anomaly detection.
3.3 Bridging Theoretical Modeling and Practical Anomaly Detection with Relative Anomaly Distribution Index
3.3.1 Probability Distribution Models of Normal and Anomalous Pixel Scores
To facilitate the mathematical treatment later, we made the following two assumptions, which were verified in subsequent experiments, please see  Sec. 4.
Distribution of Normal Pixel Prediction Scores
At an Aberrance Retention Quotient of $\textit{ARQ}=\theta$, the prediction scores of normal pixels $S_{n}$ follow a normal distribution dependent on $\theta$:
$S_{n}\sim N(\mu_{n}(\theta),\sigma_{n}(\theta)^{2}),$ (3)
where the mean $\mu_{n}(\theta)$ and variance $\sigma_{n}(\theta)^{2}$ vary with ARQ. As the ARQ increases, the model’s memory of normal samples is enhanced, potentially resulting in:
  • The mean $\mu_{n}(\theta)$ becoming closer to the average value of the training data.
  • The variance $\sigma_{n}(\theta)^{2}$ decreasing due to the model’s predictions on normal samples becoming more stable.
The variance of normal pixel prediction scores decreases exponentially with increasing ARQ, expressed as:
$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta},$ (4)
where $\sigma_{n0}$ is the initial variance (without overfitting), and $k$ is a positive constant representing the rate at which variance decreases. This relationship shows that as the model overfits, its prediction on normal samples becomes more stable.
Distribution of Anomalous Pixel Prediction Scores
The prediction scores of anomalous pixels $S_{a}$ follow a fixed normal distribution that does not change with ARQ:
$S_{a}\sim N(\mu_{a},\sigma_{a}^{2}),$ (5)
since anomalous pixels are not present during training, and the model cannot memorize them.
3.3.2 Introduction of RADI
After these assumptions, we introduce the Relative Anomaly Distribution Index, called RADI, a novel metric designed to bridge theoretical modeling and practical anomaly detection. Unlike traditional metrics such as AUROC, which are often used to evaluate model performance in anomaly detection tasks, RADI provides an indirect yet insightful measure of the model’s capacity to differentiate between normal and anomalous samples. This makes it particularly valuable for theoretical analyses, while also retaining practical utility in assessing model effectiveness.
The RADI quantifies the overlap between normal and anomalous score distributions, akin to the Wilcoxon rank-sum [24] and Mann-Whitney U tests [19]. RADI is calculated using the cumulative distribution function (CDF) of the score distributions:
[…]
where $S_{a}$ and $S_{n}$ represent the scores of anomalous and normal pixels, respectively.
It might seem counterintuitive why using RADI as an indicator would enhance model performance, as one might think that maximizing this value would lead to the model predicting everything as anomalous. However, RADI itself is not directly used for model optimization; rather, it serves as a metric to reflect the optimization outcome and help identify the extent to which normal and anomalous distributions are distinguishable. The goal is to ensure that anomalies receive consistently higher scores compared to normal samples, which allows the model to achieve better discrimination.
Moreover, the relationship between overfitting and variance is controlled by  Eq. 8, which ensures that the variance of normal pixel predictions ($\sigma_{n}(\theta)$) does not decrease to zero due to the noise term $\sigma_{\text{noise}}(\theta)$. This prevents the variance from collapsing entirely, maintaining a distinction between normal and anomalous scores. Thus, RADI can increase without the model predicting all samples as anomalous, as variability in normal predictions remains above a non-zero threshold, sustaining a balance between sensitivity to anomalies and generalization.
RADI provides a probabilistic assessment of whether anomalous scores exceed normal scores, offering insights into model discrimination. While similar to AUROC in reflecting a model’s ability to distinguish between normal and anomalous samples, RADI calculates this directly through distribution comparison rather than at multiple thresholds. It is particularly useful for quick theoretical assessments, whereas AUROC is better suited for comprehensive performance evaluations across thresholds.
Overall, RADI complements AUROC by providing a simplified measure of the model’s discriminative power, effectively reflecting its trend without requiring threshold-based evaluations.
3.3.3 Theoretical Derivation of Optimal Overfitting Conditions
For our case, where both $S_{a}$ and $S_{n}$ follow normal distributions, RADI can be further expressed in a simplified analytical form:
$\textit{RADI}(\theta)=\Phi\left(\frac{\mu_{a}-\mu_{n}}{\sqrt{\sigma_{n}(\theta)^{2}+\sigma_{a}^{2}}}\right),$ (7)
where $\Phi$ is the CDF of the standard normal distribution. In this case, $\mu_{a}$ and $\mu_{n}$ are the means of the anomaly and normal pixel score distributions, respectively, while $\sigma_{a}$ and $\sigma_{n}(\theta)$ are their respective standard deviations, with $\theta$ representing the Aberrance Retention Quotient.
Eq. 7 highlights the dependence of RADI on the statistical properties of the score distributions. As $\theta$ increases, the standard deviation $\sigma_{n}(\theta)$ of the normal scores will decrease due to the model’s increasing ability to memorize normal patterns, which results in a more stable distribution for normal scores. This reduction in variance directly affects the value of RADI, increasing the probability that the anomalous scores are distinguished from normal scores, thereby enhancing the detection capability of the model.
However, when overfitting is excessive, the model may start to memorize noise, leading to instability in prediction scores, and the variance $\sigma_{n}(\theta)$ will no longer decrease and could even increase. To describe this phenomenon, we introduce a noise term:
$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta}+\sigma_{\text{noise}}(\theta),$ (8)
where $\sigma_{\text{noise}}(\theta)$ represents the noise impact induced by overfitting and can be expressed as:
$\sigma_{\text{noise}}(\theta)=\sigma_{\text{max}}(1-e^{-h(\theta-\theta_{0})}),\quad\text{for}\;\theta>\theta_{0},$ (9)
where $\sigma_{\text{max}}$ is the maximum noise variance, $h$ controls the rate of increase of noise variance, and $\theta_{0}$ is the threshold of ARQ where noise starts to appear. When $\theta\leq\theta_{0}$, $\sigma_{\text{noise}}(\theta)=0$.
Category DSR [31] PatchCore [21] BGAD [26] SimpleNet [17] GLASS [4] RD [6] RD (+COAD) RD++ [23] RD++ (+COAD)
Object Bottle 99.6/98.8 100/98.5 100/98.9 100/98.0 100/99.3 100/98.7 100/98.8 100/98.8 100/99.6
Cable 95.3/97.7 99.8/98.4 97.9/98.0 100/97.5 99.8/98.7 95.0/97.3 97.9/97.7 99.3/98.4 99.4/99.2
Capsule 98.3/91.0 98.1/99.0 97.3/98.0 97.8/98.9 99.9/99.4 96.3/98.3 98.8/98.4 99.0/98.8 99.7/99.5
Hazelnut 97.7/99.1 100/98.7 99.3/98.5 99.8/98.1 100/99.4 99.9/98.9 100/99.1 100/99.2 100/99.4
Metal Nut 99.1/94.1 100/98.3 99.3/97.7 100/98.8 100/99.4 100/97.3 100/97.7 100/98.1 100/99.6
Pill 98.9/94.2 96.4/97.8 98.8/98.0 98.6/98.6 99.3/99.4 96.6/98.2 97.9/98.4 98.4/98.3 99.5/99.0
Screw 95.9/98.1 98.4/99.5 92.3/99.2 98.7/99.2 100/99.5 97.0/99.5 99.5/99.6 98.9/99.7 99.9/99.6
Toothbrush 100/99.5 100/98.6 86.9/98.7 100/98.5 100/99.3 99.5/98.9 100/99.1 100/99.1 100/99.4
Transistor 96.3/80.3 99.9/96.1 99.7/93.9 100/97.0 99.9/97.6 96.7/92.4 98.9/93.4 98.5/94.3 99.7/98.9
Zipper 98.5/98.4 99.4/98.9 97.8/98.7 99.9/98.9 100/99.6 98.5/98.2 99.1/98.7 98.6/98.8 99.8/99.5
Texture Carpet 99.6/96.0 98.6/99.1 99.8/99.4 99.7/98.4 99.8/99.6 98.9/98.9 99.8/99.2 100/99.2 99.8/99.5
Grid 100/99.6 97.7/98.8 99.1/99.4 99.9/98.5 100/99.4 100/99.1 100/99.3 100/99.3 100/99.5
Leather 99.3/99.5 100/99.3 100/99.7 100/99.2 100/99.8 100/99.4 100/99.5 100/99.5 100/99.9
Tile 100/98.6 98.8/95.7 100/96.7 98.7/97.7 100/99.7 99.3/95.6 99.9/95.7 99.7/96.6 100/98.8
Wood 94.7/91.5 99.1/95.0 99.5/97.0 99.5/94.4 99.9/98.8 99.2/95.3 99.6/95.5 99.3/95.8 99.9/98.8
Avg. 98.2/95.8 99.1/98.1 97.9/98.2 99.5/98.1 99.9/99.3 98.5/97.7 99.4/98.0 99.4/98.3 99.9/99.4
Table 1: Performance comparison across different methods in one-class tasks. Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method. The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
To find the optimal Aberrance Retention Quotient of $\textit{ARQ}=\theta^{*}$ that maximizes RADI, we calculate the derivative of $\text{AUROC}_{\text{pixel}}(\theta)$ and set it to zero.
The derivative is as below and we set the derivative to zero:
$\frac{d\textit{RADI}(\theta)}{d\theta}=\phi(z(\theta))\cdot\frac{dz(\theta)}{d\theta}=0,$ (10)
where $\phi(z)$ is the probability density function of the standard normal distribution.
$\frac{dz(\theta)}{d\theta}=-\frac{(\mu_{a}-\mu_{n})\cdot\sigma_{n}(\theta)\cdot\frac{d\sigma_{n}(\theta)}{d\theta}}{(\sigma_{n}(\theta)^{2}+\sigma_{a}^{2})^{3/2}}=0,$ (11)
This implies either $\sigma_{n}(\theta)=0$ (unrealistic as variance cannot be zero) or $\frac{d\sigma_{n}(\theta)}{d\theta}=0$. Thus, solving yields:
$\theta^{*}=\frac{\ln(k\sigma_{n0})-\ln(h\sigma_{\text{max}})+h\theta_{0}}{k-h}.$ (12)
3.3.4 Dual Control Mechanism and Freeze Command Signals
We define Aberrance Retention Quotient of $\textit{ARQ}=\theta$, adjusted to optimize the pixel-level AUROC. The mathematical relationship between ARQ and AUROC is established by monitoring the gradient, We call it Gradient-Guided Overfitting Control.
$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0.$ (13)
We integrate $ARQ_{optimal}$ in  Eq. 2 and Gradient-Guided Overfitting Control in  Eq. 13 and call this combined control mechanism the Dual Control Mechanism:
$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0,\quad\text{for}\;\theta\in\textit{$ARQ_{optimal}$}.$ (14)
If Dual Control Mechanism fails, the Dual Control Mechanism triggers Freeze Command Signals, which prompt whether freeze certain network layers, beginning with lower-level feature extraction layers and moving upwards.
The rationale for progressively freezing certain layers while overfitting the other parts lies in the architecture’s functional partitioning. Lower layers are responsible for extracting fundamental features, which are common across both normal and anomalous data. Freezing these layers early preserves their ability to extract generalizable features, preventing degradation due to overfitting. Conversely, keeping deeper layers unfrozen during the overfitting stage enables the model to specialize in identifying unique, high-level features that are more discriminative between normal and anomalous samples. This approach helps maintain generalization while enhancing sensitivity to rare patterns, ultimately boosting anomaly detection performance without succumbing to the pitfalls of overfitting the entire model.
4 Experiments
Category US [2] PSVDD [27] PaDiM [5] MKD [22] DRAEM [30] UniAD [28] UniAD (+COAD) DiAD [11] DiAD (+COAD)
Object Bottle 84.0/67.9 85.5/86.7 97.9/96.1 98.7/91.8 97.5/87.6 99.7/98.1 100/98.0 99.7/98.4 100/98.8
Cable 60.0/78.3 64.4/62.2 70.9/81.0 78.2/89.3 57.8/71.3 95.2/97.3 97.9/98.8 94.8/96.8 97.5/97.7
Capsule 57.6/85.5 61.3/83.1 73.4/96.9 68.3/88.3 65.3/50.5 86.9/98.5 92.3/98.6 89.0/97.1 98.0/98.8
Hazelnut 95.8/93.7 83.9/97.4 85.5/96.3 97.1/91.2 93.7/96.9 99.8/98.1 99.9/98.1 99.5/98.3 100/99.1
Metal Nut 62.7/76.6 80.9/96.0 88.0/84.8 64.9/64.2 72.8/62.2 99.2/94.8 99.3/93.2 99.1/97.3 100/97.7
Pill 56.1/80.3 89.4/96.5 68.8/87.7 79.7/69.7 82.2/94.4 93.7/95.0 94.4/98.4 95.7/95.7 97.4/98.4
Screw 66.9/90.8 80.9/74.3 56.9/94.1 75.6/92.1 92.0/95.5 87.5/98.3 90.6/98.6 90.7/97.9 99.5/99.6
Toothbrush 57.8/86.9 99.4/98.0 95.3/95.6 75.3/88.9 90.6/97.7 94.2/98.4 94.7/98.6 99.7/99.0 100/99.1
Transistor 61.0/68.3 77.5/78.5 86.6/92.3 73.4/71.7 74.8/64.5 99.8/97.9 99.8/98.4 99.8/95.1 99.8/93.4
Zipper 78.6/84.2 77.8/95.1 79.7/94.8 87.4/86.1 98.8/98.3 95.8/96.8 97.6/98.1 95.1/96.2 98.3/98.7
Texture Carpet 86.6/88.7 63.3/78.6 93.8/97.6 69.8/95.5 98.0/98.6 99.8/98.5 100/98.8 99.4/98.6 99.6/99.2
Grid 69.2/64.5 66.0/70.8 73.9/71.0 83.8/82.3 99.3/98.7 98.2/96.5 99.1/98.4 98.5/96.6 100/99.3
Leather 97.2/95.4 60.8/93.5 99.9/84.8 93.6/96.7 98.7/97.3 100/98.8 100/98.8 99.8/98.8 100/99.5
Tile 93.7/82.7 88.3/92.1 93.3/80.5 89.5/85.3 99.8/98.0 99.3/91.8 99.3/98.6 96.8/92.4 99.3/95.7
Wood 90.6/83.3 72.1/80.7 98.4/89.1 93.4/80.5 99.8/96.0 98.6/93.2 98.5/96.1 99.7/93.3 98.7/95.5
Avg. 74.5/81.8 76.8/85.6 84.2/89.5 81.9/84.9 88.1/87.2 96.5/96.8 97.6/98.0 97.2/96.8 99.1/98.0
Table 2: Performance comparison across different methods in multi-class tasks. Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method. The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
4.1 Dataset
We evaluate our method using the MVTec AD dataset [1], which is a comprehensive, real-world dataset for unsupervised anomaly detection tasks. The MVTec AD dataset includes 15 categories, comprising 3629 images for training and 1725 images for testing. The dataset provides pixel-level annotations for each anomaly, making it suitable for evaluating both image-level classification and pixel-level segmentation tasks.
4.2 Comparison with SOTAs
In this section, we compare the performance of our proposed method with several SOTA anomaly detection methods in both one-class and multi-class tasks.
Using the optimal ARQ value $\theta$ derived from the $ARQ_{o}ptimal$ as indicated by  Eq. 12, we set ARQ to 0.006 for one-class tasks and 0.06 for multi-class tasks. $ARQ_{optimal}$ is defined as $\theta\in[0.001,0.011]$ for one-class tasks and $\theta\in[0.01,0.11]$ for multi-class tasks, where the balance between overfitting and generalization is optimal for anomaly detection. We saved model states within these intervals and evaluated their performance across respective benchmarks to ensure robust detection.
Furthermore, during model training, we adopted a Dual Control Mechanism for Optimal Detection Performance approach in our framework COAD, as detailed in  Eq. 14, and when Dual Control Mechanism fails, it triggers Freeze Command Signals, prompting the selective freezing of certain network layers in the traditional framework.
Our method builds upon existing anomaly detection frameworks. Initially, we perform standard training on the original model. Following this, we transition into the controllable overfitting stage, during which the model is further trained using our COAD framework. This stage utilizes Dual Control Mechanism in  Eq. 14. All pseudo-anomaly generation in this stage is performed using Gaussian noise to introduce controlled variability, which aids in refining the model’s sensitivity to anomalies. Additionally, the learning rate is reduced to $\frac{1}{10}$ of that used in the normal training stage, allowing the model to make finer adjustments.
After training, the inference of the model proceeds in the same manner as the baseline anomaly detection frameworks, ensuring that the enhancements provided by COAD integrate seamlessly.
The results of our one-class and multi-class experiments are summarized in  Tabs. 1 and 2. It can be observed that our enhanced method outperforms the original methods, achieving new SOTA results. The enhancement in pixel-level AUROC is especially significant, which is crucial for fine-grained anomaly detection.
These results confirm that the introduction of controllable overfitting, regulated by ARQ and RADI, enhances the model’s sensitivity and stability, leading to robust anomaly detection abilities that outperform existing SOTA methods.
4.3 Ablation Study
4.3.1 Comparisons with Our Basic Frameworks
RD RD++ UniAD DiAD
Baseline 98.5 99.4 96.5 97.2
+COAD 99.4(+0.9) 99.9(+0.5) 97.6(+1.1) 99.1(+1.9)
Table 3: Image-Level AUROC comparison between the basic and enhanced versions of different frameworks.
RD RD++ UniAD DiAD
Baseline 97.7 98.3 96.8 96.8
+COAD 98.0(+0.3) 99.4(+1.1) 98.0(+1.2) 98.0(+1.2)
Table 4: Pixel-Level AUROC comparison between the basic and enhanced versions of different frameworks.
We conduct an ablation study comparing our enhanced framework with its corresponding basic versions in  Tabs. 3 and 4. The results demonstrate consistent and notable improvements in both image-level and pixel-level AUROC across all evaluated frameworks.
In particular, the performance gains observed in image-level anomaly detection emphasize the strength of our COAD, which effectively captures subtle differences between normal and anomalous regions. Specifically, our enhancements yielded an increase of up to 1.9 in the DiAD framework, as seen in image-level AUROC, and similar significant improvements in other frameworks.
For pixel-level AUROC, a similar pattern of improvement was observed, with gains of up to 1.2 seen in multiple frameworks, as shown in Tab. 4. This underscores the strength of our COAD, which not only optimizes image-level detection but also provides enhanced granularity in distinguishing pixel-level abnormalities.
Furthermore, the versatility of our method is evident in its applicability to different anomaly detection paradigms. Whether in one-class or multi-class anomaly detection tasks, or even in settings employing diffusion-based models versus non-diffusion frameworks, our method has consistently demonstrated superior performance, showcasing our method’s robustness and generalizability.
4.3.2 Comparisons with Distinct ARQ
Model 0–0.001 0.001–0.011 0.011–0.1
RD 98.8 / 97.8 99.4 / 98.0 99.6 / 97.6
RD++ 99.7 / 98.8 99.9 / 99.4 -
Table 5: Comparison of performance across distinct ARQ ranges for different models in one-class tasks.
Model 0–0.01 0.01–0.11 0.11–1
UniAD 97.0 / 96.8 97.6 / 98.0 98.1 / 96.9
DiAD 97.3 / 96.6 99.1 / 98.0 99.2 / 96.5
Table 6: Comparison of performance across distinct ARQ ranges for different models in multi-class tasks.
In this section, we compare the performance of different models under distinct ARQ ranges to validate the impact of controllable overfitting on anomaly detection. We evaluated methods such as RD, RD++, UniAD, and DiAD across various ARQ levels. As  Tabs. 5 and 6 show, the results demonstrate that, by appropriately regulating the ARQ, which is in $ARQ_{optimal}$, model detection performance can be significantly enhanced. It is evident that our controllable overfitting strategy not only yields consistent improvements across different models but also helps in striking the optimal balance between overfitting and generalization.
Overall, these results demonstrate that our controllable overfitting strategy consistently yields significant performance gains across various models and tasks. By leveraging the ARQ within the identified optimal range, we successfully enhance the models’ discriminative abilities while maintaining a balanced relationship between overfitting and generalization.
4.4 Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
ARQ $\mu_{n}$ $\theta_{n}$ $\mu_{a}$ $\theta_{a}$ $TVD_{n}$ $TVD_{a}$
0 105.6 65.5 99.3 45.0 0.20 0.08
0.006 105.3 65.5 98.9 44.9 0.20 0.08
0.015 105.6 65.5 99.0 45.1 0.20 0.08
0.045 105.8 65.4 99.3 45.1 0.20 0.08
0.050 105.6 65.3 99.1 45.0 0.21 0.09
0.060 105.6 65.2 99.0 45.0 0.20 0.09
0.082 105.5 65.0 99.0 45.0 0.19 0.08
Table 7: Partial Results of Validation Metrics for Assumptions on Normal and Anomalous Pixel Score Distributions.
In this section, we firstly validate the assumptions introduced in Sec. 3.3.1 by employing the Total Variation Distance (TVD) metric. We leveraging multiple models at different stages of overfitting, characterized by distinct ARQ values. For each model state, corresponding to different ARQ = $\theta$, we calculate the prediction scores for both normal and anomalous pixels across the dataset. Subsequently, we evaluate their proximity to a Gaussian distribution using TVD, ensuring a robust assessment of the normality of these distributions. The partial results of this validation are summarized in Tab. 7. Moreover, we model the variance of the normal pixel prediction scores $(\sigma_{n}(\theta))$ as an exponential function of the ARQ = $\theta$, given by Eq. 4.
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we employ the TVD metric to quantitatively compare the distribution of real anomaly scores with that of Gaussian-distributed pseudo-anomalies. We can see  Tab. 7, the resulting TVD value was about 0.08, suggesting that the Gaussian-distributed pseudo-anomalies exhibit a significant overlap with the actual anomalies. This supports the assumptions that Gaussian noise can be effectively used to preliminarily simulate anomalous behavior in training models for anomaly detection, which is usually used by GLASS[4], SimpleNet[17], RealNet[32],and DDPM [12].
5 Conclusion
In this work, we reenvisioned overfitting as a controllable and transformative mechanism, using COAD to enhance model capabilities beyond conventional boundaries. We introduced ARQ to precisely regulate overfitting, as well as RADI, which leverages $ARQ_{optimal}$ to provide a more versatile metric compared to AUROC-pixel, facilitating both theoretical modeling and Dual Control Mechanism. By repurposing overfitting as a generalizable module, our approach not only dismantles its demonization but also enables us to achieve SOTA results in both one-class and multi-class anomaly detection tasks. Furthermore, we provide a robust theoretical foundation for employing Gaussian noise as a preliminary pseudo-anomaly generator, extending the applicability of our COAD.
Supplementary Material
6 Dual Control Mechanism Design Concept
Overfitting has traditionally been viewed as a problem that hampers generalization—where the model memorizes noise and specific details of the training data, leading to poor performance on unseen examples. However, in the context of anomaly detection, we posit that overfitting can be harnessed as a feature, not a flaw. The key lies in the concept of controllable overfitting—managing overfitting in a way that makes the model more sensitive to subtle differences between normal and anomalous data.
In practical terms, overfitting can amplify the sensitivity of the model to slight deviations, which is particularly useful in anomaly detection scenarios, where anomalies are often defined by subtle discrepancies from the norm. Our approach aims to leverage this characteristic to enhance detection capability while avoiding the negative consequences of overfitting, such as loss of generalization. This balancing act necessitates a carefully designed mechanism, which we call the Dual Control Mechanism in  Eq. 14.
The Dual Control Mechanism is comprised of two complementary components: $ARQ_{optimal}$ in  Eq. 2 and Gradient-Guided Overfitting Control in  Eq. 13. These components work together to ensure that overfitting remains within beneficial bounds, ultimately enhancing anomaly detection without sacrificing the ability to generalize effectively.
Aberrance Retention Quotient
The ARQ quantifies the degree of overfitting by capturing the relationship between the model’s fit to the training data and its potential for retaining useful deviations that indicate anomalies. We define an optimal interval for ARQ, denoted as ARQoptimal, which represents the range where overfitting is beneficial—significantly enhancing sensitivity to anomalies without compromising generalization.
During training, ARQ is continuously monitored to ensure it remains within this optimal interval. If ARQ moves outside the defined range, corrective actions are taken to bring it back. The optimal interval allows the model to exploit overfitting in a targeted manner, enhancing the model’s sensitivity towards subtle anomalies while minimizing the risk of memorizing noise.
Gradient-Guided Overfitting Control
Alongside ARQoptimal, the Gradient-Guided Overfitting Control is employed to supervise how overfitting impacts anomaly detection performance directly. This component utilizes the Relative Anomaly Distribution Index (RADI) to quantify the overlap between normal and anomalous score distributions. By keeping track of the gradient of RADI, we can ensure that the anomaly detection performance is on an upward trajectory during overfitting.
Specifically, we require that the gradient of RADI with respect to ARQ is non-negative, which indicates that the model’s ability to distinguish anomalies from normal data is improving as overfitting increases. If the gradient condition is violated, indicating that overfitting is leading to a decline in detection performance, additional measures are triggered.
Freeze Command Signals for Progressive Layer Freezing
In situations where the Dual Control Mechanism fails—meaning both the ARQ exceeds the optimal range and the Gradient-Guided Overfitting Control shows deteriorating performance—we initiate Freeze Command Signals in. These signals are the model’s way of mitigating excessive overfitting by progressively freezing network layers.
Freeze Command Signals begin by selectively freezing the lower-level feature extraction layers, which are primarily responsible for capturing fundamental visual features such as edges, textures, and basic shapes. These foundational features are generally invariant across different images and therefore require less flexibility once they are well-trained. By freezing these lower-level layers, we stabilize the feature extraction process, reducing the risk of the model overfitting to noise and minute variations in the training set. As training progresses and overfitting persists beyond the optimal range, additional higher-level layers are frozen progressively in a staged manner. This gradual freezing strategy ensures that while the foundational layers remain stable, the more abstract, task-specific layers retain flexibility long enough to refine their understanding of anomalies.
By progressively freezing layers, the model maintains stable low-level feature extraction while allowing the higher-level layers to refine their understanding of anomalies. This approach ensures that the model retains a balance between maintaining its basic feature extraction capabilities and enhancing its discriminative power for anomaly detection.
The Dual Control Mechanism thus provides a comprehensive way to manage overfitting during training. By monitoring both ARQ and the RADI gradient, and employing Freeze Command Signals when necessary, our framework transforms overfitting from a challenge into an asset—amplifying the model’s ability to detect anomalies while maintaining robustness and generalization.
7 Main Functional Pseudocode
Algorithm 1 Training Stage with Gaussian Noise, Teacher-Student Reconstruction, and Overfitting Control
0:   Dataset $\mathcal{D}$, Initial Model $M$, Learning Rate $\alpha$, Optimal ARQ Range $\theta_{optimal}$, Noise Standard Deviation $\sigma_{noise}$, Epochs $N_{standard}$, Overfit Epochs $N_{overfit}$
1:   Step 1: Standard Training Stage
2:   for epoch = 1 to $N_{standard}$ do
3:      for batch in $\mathcal{D}_{train}$ do
4:         Generate pseudo-anomalous data $\mathcal{D}_{anomaly}$ using Gaussian noise.
5:         Generate feature maps $F_{T}$ using the Teacher Model $M_{T}$ for training samples.
6:         Generate reconstructed feature maps $F_{S}$ using the Student Model $M_{S}$.
7:         Compute $L_{reconstruction}$ between $F_{T}$ and $F_{S}$ and other losses.
8:         Update Models
9:      end for
10:   end for
11:   Step 2: Controllable Overfitting Stage
12:   Set learning rate $\alpha_{overfit}=\alpha/10$
13:   Initialize Freeze Counter $C=0$
14:   for epoch = 1 to $N_{overfit}$ do
15:      for batch in $\mathcal{D}_{train}$ do
16:         Generate pseudo-anomalous data $\mathcal{D}_{anomaly}$ using Gaussian noise.
17:         Generate feature maps $F_{T}$ using the Teacher Model $M_{T}$ for both normal and pseudo-anomalous samples.
18:         Generate reconstructed feature maps $F_{S}$ using the Student Model $M_{S}$.
19:         Calculate ARQ
20:         Compute RADI based on score distributions
21:         Dual Control Mechanism:
22:         if $\theta$ exceeds $\theta_{optimal}$ or $\frac{dRADI(\theta)}{d\theta}<0$ then
23:            Increment Freeze Counter $C$
24:            if $C>C_{t}hr$ then
25:               Issue Freeze Command Signals
26:               Selective Freezing
27:            end if
28:         else
29:            Reset Freeze Counter $C$
30:         end if
31:         Compute $L_{reconstruction}$ between $F_{T}$ and $F_{S}$ and other losses.
32:         Update Models
33:      end for
34:   end for
Algorithm 2 Selective Freezing
0:   Student Model $M_{S}$, Freeze Counter $C$, Threshold $C_{max}$, Layer List $L$, ARQ $\theta$, Optimal ARQ Range $\theta_{optimal}$
1:   Initialize $C=0$Freeze Counter initialized to zero}
2:   for epoch = 1 to $N_{overfit}$ do
3:      for batch in training data do
4:         Perform forward pass and compute ARQ $\theta$
5:         if $\theta$ exceeds $\theta_{optimal}$ or $\frac{dRADI(\theta)}{d\theta}<0$ then
6:            Increment Freeze Counter: $C=C+1$
7:         else
8:            Reset Freeze Counter: $C=0$
9:         end if
10:         if $C>C_{thr}$ then
11:            Issue Freeze Command Signal to initiate layer freezing
12:            Layer Freezing Process:
13:            for layer $l$ in $L$ (ordered from lowest-level to highest-level) do
14:               if layer $l$ is not frozen then
15:                  Freeze Procedure for Current Layer:
16:                  $\quad\forall p\in\text{parameters of layer $l$}$
17:                  $\quad\quad p.requires\_grad=\textbf{False}$Stop gradient updates for layer $l$}
18:                  Update training status: mark layer $l$ as frozen
19:                  Apply Forward Freezing Effects:
20:                        Notify subsequent layers in network to adjust activations {Maintain balance during backpropagation}
21:                        Update batch normalization layers if necessary {Handle training vs. inference mode}
22:                  Reset Freeze Counter: $C=0$Reset counter after successfully freezing a layer}
23:                  Add layer $l$ to the Frozen Layer List $F$Keep track of all frozen layers}
24:                  Break from the loop {Freeze only one layer in each iteration to avoid rapid performance drops}
25:               else
26:                  Continue to next layer {Skip if layer is already frozen}
27:               end if
28:            end for
29:         end if
30:         Continue Training:
31:         Perform backpropagation and update weights accordingly
32:      end for
33:   end for
Algorithm 3 Inference Stage
0:   Trained Teacher Model $M_{\text{teacher}}$, Trained Student Model $M_{\text{student}}$, Test Dataset $\mathcal{D}_{test}$, Optional Pseudo Anomaly Filter $F$
1:   Initialization: Set models in evaluation mode.
2:   for test sample $x$ in $\mathcal{D}_{test}$ do
3:      Input Feature Extraction:
4:       Extract features using teacher model: $f_{\text{teacher}}\leftarrow M_{\text{teacher}}(x)$
5:      Student Model Inference:
6:       Extract reconstructed features using student model: $\hat{f}_{\text{student}}\leftarrow M_{\text{student}}(f_{\text{filtered}})$
7:      Reconstruction Computation:
8:       Compute reconstruction error to classify sample: $e(x)=\text{Loss}(x,\hat{f}_{\text{student}})$
9:      Prediction Generation:
10:      if $e(x)>\text{Threshold}$ then
11:         Classify as Anomalous
12:      else
13:         Classify as Normal
14:      end if
15:      Store Results: Save prediction results for sample $x$.
16:   end for
17:   Output: Anomaly classification for all test samples in $\mathcal{D}_{test}$
Algorithm 4 Calculation of Aberrance Retention Quotient (ARQ)
0:   Predicted values $\hat{y}_{i}$, Ground truth values $y_{i}$, Total instances $N$
1:   Initialize ARQ value: $\textit{ARQ}\leftarrow 0$
2:   Compute Numerator:
3:    $Numerator\leftarrow\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|$
4:   Compute Denominator:
5:    $Denominator\leftarrow\sum_{i=1}^{N}y_{i}$
6:   Calculate ARQ:
7:    $\textit{ARQ}\leftarrow\frac{Numerator}{Denominator}$
8:   Output: Final ARQ value
Algorithm 5 Calculation of Relative Anomaly Distribution Index (RADI)
0:   Scores of anomalous pixels $S_{a}$, Scores of normal pixels $S_{n}$
1:   Initialize RADI value: $\textit{RADI}\leftarrow 0$
2:   Compute RADI using Integral:
[…]
4:   Output: Final RADI value
8 Extended Discussion of Gaussian Noise Validation
8.1 Theoretical Rationale for Gaussian Noise
In this work, we chose to use Gaussian noise as a preliminarily pseudo-anomaly generator due to its statistical properties and simplicity. Gaussian noise, characterized by its mean and variance, serves as an ideal candidate for generating random deviations that can mimic unexpected patterns in normal data.
From a theoretical standpoint, Gaussian noise exhibits a distribution that is both isotropic and well-defined in high-dimensional space, making it suitable for approximating unstructured, irregular anomalies. The Central Limit Theorem (CLT) supports the use of Gaussian distributions as approximations for many naturally occurring random processes. In particular, when adding minor perturbations to an image, we aim to simulate unforeseen deviations from the underlying structure of normal samples, which Gaussian noise effectively captures.
Additionally, Gaussian noise serves as a zero-mean stochastic process, allowing us to avoid introducing any specific directional bias, which could potentially interfere with the model’s capacity to learn anomalous versus normal regions. By employing Gaussian noise, we ensure that our model does not overfit to particular anomaly patterns but instead gains a generalized sensitivity to any divergence from the normal data distribution. This universality underpins our approach to controllable overfitting within the COAD framework.
8.2 Validation Through Total Variation Distance (TVD)
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we conducted an empirical analysis using the Total Variation Distance (TVD) metric. TVD is a standard measure used in probability theory to quantify the dissimilarity between two distributions. We calculated TVD to evaluate how closely Gaussian noise could mimic true anomalies in both spatial and intensity domains.
The calculation of TVD involves estimating the difference between the empirical distributions of Gaussian noise and the true normal and anomaly pixel intensities:
$D_{\text{TVD}}(P,Q)=\frac{1}{2}\sum_{x}|P(x)-Q(x)|,$ (15)
where $P$ represents the empirical distribution of the true anomalies, and $Q$ represents the empirical distribution of Gaussian noise. TVD measures the extent of overlap between the two distributions, where a smaller TVD implies a closer approximation between the pseudo-anomalies and real anomalies.
We conducted experiments on our dataset using both generated Gaussian noise and real anomalies. presents a comparison of the TVD values for different classes. The results indicate that the Gaussian distribution closely matches the distribution of real normal and anomalies across a range of classes. Specifically, it can be seen in  Sec. 4.4.
8.3 Experimental Results and Qualitative Study
The experimental results demonstrate that the model trained using Gaussian noise exhibits a better anomaly detection performance to models trained with other pseudo-anomaly generator.
To further validate the results, we present detailed histograms for the anomaly detection scores obtained across 15 individual categories as well as the overall dataset. These histograms were generated under the condition where the Anomaly Rate Quotient (ARQ) is set to 0.006, using RD++ as our baseline framework. Each histogram in  Figs. 3 and 4 provides a clear visualization of the distribution of detection scores, overlaid with a Gaussian fitting curve for enhanced interpretability. This fitting curve highlights the distribution, allowing for a more effective comparison of the model’s performance across different categories and the overall dataset, further solidifying the reliability of our evaluation metrics.
(a) Bottle (Anomaly)
(b) Cable (Anomaly)
(c) Capsule (Anomaly)
(d) Carpet (Anomaly)
(e) Grid (Anomaly)
(f) Hazelnut (Anomaly)
(g) Leather (Anomaly)
(h) Metal Nut (Anomaly)
(i) Pill (Anomaly)
(j) Screw (Anomaly)
(k) Tile (Anomaly)
(l) Toothbrush (Anomaly)
(m) Transistor (Anomaly)
(n) Wood (Anomaly)
(o) Zipper (Anomaly)
Figure 3: Histograms for anomaly scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
(a) Bottle (Normal)
(b) Cable (Normal)
(c) Capsule (Normal)
(d) Carpet (Normal)
(e) Grid (Normal)
(f) Hazelnut (Normal)
(g) Leather (Normal)
(h) Metal Nut (Normal)
(i) Pill (Normal)
(j) Screw (Normal)
(k) Tile (Normal)
(l) Toothbrush (Normal)
(m) Transistor (Normal)
(n) Wood (Normal)
(o) Zipper (Normal)
Figure 4: Histograms for normal scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
8.4 Future Directions with GLASS-like Pseudo-Anomaly Generators
While Gaussian noise has proven effective as a preliminary pseudo-anomaly generator, future work could explore more advanced methods inspired by the GLASS approach [4]. The GLASS method introduces a structured way to generate pseudo-anomalies by leveraging geometric transformations and localized pattern synthesis, providing greater control over anomaly characteristics.
By incorporating concepts from GLASS, we envision a Gaussian Noise-Enhanced GLASS-like Generator, where Gaussian noise could act as the foundational perturbation, and additional constraints or transformations could be applied to shape the pseudo-anomalies in a contextually relevant manner. For instance, augmenting Gaussian noise with spatial or intensity correlations reflective of real anomalies could enhance its realism and improve model robustness in detecting diverse anomalies.
Such an approach would align well with our COAD framework, enabling a more versatile training process that accounts for both unstructured and structured anomaly patterns. By systematically integrating these future developments, we aim to refine our anomaly detection strategy, bridging the gap between theoretical advancements and practical applications.
9 Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
To further validate the effectiveness of our methods, we provide a qualitative comparison of anomaly detection results across 15 distinct categories using baselines such as RD, RD++, UniAD, DiAD, and the proposed COAD framework in  Figs. 5, 6, 7 and 8. Each subfigure demonstrates: (from left to right) the original image, the ground truth anomaly mask, the result from the baseline method, and the result after applying COAD.
The comparison shows that COAD effectively reduces both false positive rates (FPR) and false negative rates (FNR), accurately capturing anomaly regions while minimizing misclassifications. These results are consistent with the improvements reflected in the earlier quantitative data in  Tabs. 3 and 4, where COAD consistently outperforms its counterparts in image-level and pixel-level AUROC metrics, highlighting COAD’s capability to outperform the baselines.
By leveraging controllable overfitting, COAD enhances the model’s sensitivity to true anomalies while maintaining its generalization capabilities. The reduction in FPR minimizes the misclassification of normal regions as anomalies, while the decreased FNR highlights the model’s ability to capture subtle anomalies that baseline methods often miss. These qualitative and quantitative improvements collectively demonstrate the robustness and effectiveness of COAD in advancing anomaly detection tasks. Additionally, the sharper distinction between normal and anomalous regions demonstrates the framework’s capacity to refine boundary precision, further enhancing detection reliability.
(a) Bottle
(b) Cable
(c) Capsule
(d) Carpet
(e) Grid
(f) Hazelnut
(g) Leather
(h) Metalnut
(i) Pill
(j) Screw
(k) Tile
(l) Toothbrush
(m) Transistor
(n) Wood
(o) Zipper
Figure 5: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using the Baseline Method RD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
(a) Bottle
(b) Cable
(c) Capsule
(d) Carpet
(e) Grid
(f) Hazelnut
(g) Leather
(h) Metalnut
(i) Pill
(j) Screw
(k) Tile
(l) Toothbrush
(m) Transistor
(n) Wood
(o) Zipper
Figure 6: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method RD++ and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
(a) Bottle
(b) Cable
(c) Capsule
(d) Carpet
(e) Grid
(f) Hazelnut
(g) Leather
(h) Metalnut
(i) Pill
(j) Screw
(k) Tile
(l) Toothbrush
(m) Transistor
(n) Wood
(o) Zipper
Figure 7: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method UniAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
(a) Bottle
(b) Cable
(c) Capsule
(d) Carpet
(e) Grid
(f) Hazelnut
(g) Leather
(h) Metalnut
(i) Pill
(j) Screw
(k) Tile
(l) Toothbrush
(m) Transistor
(n) Wood
(o) Zipper
Figure 8: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method DiAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
10 Future Work
In future work, we plan to refine our pseudo-anomaly generators by incorporating advanced methods in  Sec. 8.4, enabling the generation of more structured and domain-relevant pseudo-anomalies.
Moreover, we intend to explore the application of the COAD framework in other domains, particularly in areas such as medical imaging, autonomous driving, and forgery detection. Forgery detection, for instance, highlights an opportunity to utilize COAD’s enhanced sensitivity to subtle anomalies, such as fine-grained pattern inconsistencies or texture deviations. Furthermore, we aim to generalize and modularize the COAD framework into a plug-and-play component that can be seamlessly integrated into similar tasks across various fields. Ultimately, we aspire to redefine the concept of overfitting itself, transforming it from a perceived limitation into a powerful tool that drives innovation.

### v2

@@
 Friend or Foe? Harnessing Controllable Overfitting for Anomaly Detection
-Abstract
-Overfitting has long been stigmatized as detrimental to model performance, especially in the context of anomaly detection.
-Our work challenges this conventional view by introducing a paradigm shift—recasting overfitting as a controllable and strategic mechanism for enhancing model discrimination capabilities.
-In this paper, we present Controllable Overfitting-based Anomaly Detection (COAD), a novel framework designed to leverage overfitting for optimized anomaly detection.
-We propose the Aberrance Retention Quotient (ARQ), a novel metric that systematically quantifies the extent of overfitting, enabling the identification of an optimal ”golden overfitting interval” $ARQ_{optimal}$.
-Within this interval, overfitting is leveraged to significantly amplify the model’s sensitivity to anomalous patterns, while preserving generalization to normal samples.
-Additionally, we present the Relative Anomaly Distribution Index (RADI), an innovative metric designed to complement AUROC-pixel by providing a more versatile and theoretically robust framework for assessing model performance.
-RADI leverages ARQ to track and evaluate how overfitting impacts anomaly detection, offering an integrated approach to understanding the relationship between overfitting dynamics and model efficacy.
-Our theoretical work also rigorously validates the use of Gaussian noise in pseudo-anomaly synthesis, providing the foundation for its broader applicability across diverse domains.
-Empirical evaluations demonstrate that our controllable overfitting method not only achieves State-Of-The-Art(SOTA) performance in both one-class and multi-class anomaly detection tasks but also redefines overfitting from a modeling challenge into a powerful tool for optimizing anomaly detection.
-1 Introduction
-In the traditional context of machine learning, overfitting has long been perceived as an undesirable phenomenon, just like  [3, 8, 10, 20], traditionally viewed as the result of a model excessively memorizing training data, leading to compromised generalization capabilities.
-However, we challenge this conventional wisdom by extending the concept beyond the domain of anomaly detection, positioning controllable overfitting as a versatile mechanism for enhancing model sensitivity across a wide range of applications.
+Abstract.
+Overfitting has traditionally been viewed as detrimental to anomaly detection, where excessive generalization often limits models’ sensitivity to subtle anomalies.
+Our work challenges this conventional view by introducing Controllable Overfitting-based Anomaly Detection (COAD), a novel framework that strategically leverages overfitting to enhance anomaly discrimination capabilities.
+We propose the Aberrance Retention Quotient (ARQ), a novel metric that systematically quantifies the extent of overfitting, enabling the identification of an optimal ”golden overfitting interval” $ARQ_{optimal}$ wherein model sensitivity to anomalies is maximized without sacrificing generalization.
+To comprehensively capture how overfitting affects detection performance, we further propose the Relative Anomaly Distribution Index (RADI), a metric superior to traditional AUROC by explicitly modeling the separation between normal and anomalous score distributions.
+Theoretically, RADI leverages ARQ to track and evaluate how overfitting impacts anomaly detection, offering an integrated approach to understanding the relationship between overfitting dynamics and model efficacy.
+We also rigorously validate the statistical efficacy of Gaussian noise as pseudo-anomaly generators, reinforcing the method’s broad applicability.
+Empirical evaluations demonstrate that our controllable overfitting method achieves State-Of-The-Art(SOTA) performance in both one-class and multi-class anomaly detection tasks, thus redefining overfitting as a powerful strategy rather than a limitation.
+Anomaly Detection, Controllable Overfitting, Aberrance Retention Quotient, Relative Anomaly Distribution Index, Gaussian Noise, Dual Control Mechanism
+ccs: Computing methodologies Scene anomaly detection
+1. Introduction
+In the traditional context of machine learning, overfitting has long been perceived as an undesirable phenomenon, just like  (Bishop, 2006; Goodfellow et al., 2016; Hastie et al., 2004; Murphy, 2012), traditionally viewed as the result of a model excessively memorizing training data, leading to compromised generalization capabilities.
+However, we challenge this conventional wisdom by extending the concept beyond the domain of anomaly detection, positioning controllable overfitting as a mechanism for enhancing model sensitivity across a wide range of applications.
 This leads us to the central question that underpins our work: can overfitting be a friend rather than a foe in the domain of anomaly detection?
@@
 (b) False Negative Cases
-Figure 1: Each subfigure presents a sequence from left to right, displaying: the original image, the ground truth, the results from other methods, like RD [6], RD++[23], UniAD [28] and DiAD [11], and our results after overfitting enhancement based on existing frameworks with COAD.
-(a) illustrates a significant reduction in the false positive rate after applying our overfitting approach.
-(b) highlights the reduction in the false negative rate, demonstrating enhanced sensitivity towards true anomalies with our method.
+Figure 1.
+Each subfigure presents a sequence from left to right, displaying: the original image, the ground truth, the results from other methods, like RD (Deng and Li, 2022), RD++(Tien et al., 2023), UniAD (You et al., 2022) and DiAD (He et al., 2023), and our results after overfitting enhancement based on existing frameworks with COAD.
+(a) illustrates a significant reduction in the false positive case after applying our overfitting approach.
+(b) highlights the reduction in the false negative case, demonstrating enhanced sensitivity towards true anomalies with our method.
 The consistent improvement across both metrics indicates the effectiveness of our methods in refining detection accuracy.
 The conventional understanding of overfitting necessitates a significant paradigm shift.
-Historically stigmatized as inherently detrimental to machine learning models, overfitting, as our research demonstrates, can under specific controlled conditions be repurposed as a powerful mechanism for enhancing model performance.
-In scenarios where training data consists solely of normal samples, we show that allowing the model to comprehensively memorize these features enhances its sensitivity to deviations—namely, anomalies—during testing.
-This reframing positions overfitting not as an inherent flaw to be avoided, but as a strategic asset to be leveraged effectively.
-Anomaly detection in visual domains is a crucial task that aims to identify patterns that do not conform to expected behavior.
+Historically stigmatized as inherently detrimental to machine learning models, overfitting, as our work demonstrates, can under specific controlled conditions be repurposed as a powerful mechanism for enhancing model performance.
+Intuitively, by enabling the model to intentionally and moderately memorize fine-grained patterns unique to normal samples, we amplify its sensitivity toward any deviations—precisely the subtle anomalies we aim to detect.
+This reframes overfitting from a phenomenon to avoid into a strategically leveraged resource.
+Anomaly detection is a crucial task that aims to identify patterns that do not conform to expected behavior.
 Traditionally, anomaly detection relies on the notion of learning the distribution of normal data and distinguishing anomalies by identifying outliers.
-Various methods have been proposed to achieve this, including reconstruction-based approaches [31, 15, 30], embedding-based models [25, 32, 13, 9], and anomaly synthesis technique [4, 28, 17, 23].
+Various methods have been proposed, including reconstruction-based approaches (Zavrtanik et al., 2022; Liang et al., 2023; Zavrtanik et al., 2021), embedding-based models (Xiao et al., 2023; Zhang et al., 2024; Lei et al., 2023; Gudovskiy et al., 2021), and anomaly synthesis technique (Chen et al., 2024; You et al., 2022; Liu et al., 2023; Tien et al., 2023).
 We advance the notion of controllable overfitting, suggesting that a calculated and moderate degree of overfitting enables the model to capture intricate and subtle characteristics of the data.
@@
 In this paper, we present Controllable Overfitting-based Anomaly Detection, called COAD.
-Our theoretical framework and empirical validation jointly support this proposition.
-We demonstrate that, as the degree of overfitting increases, the score distribution for normal samples evolves accordingly, while the score distribution for anomalous samples remains stable.
-By introducing the Aberrance Retention Quotient, called ARQ, as an overfitting parameter, coupled with a noise loss term to finely control the degree of overfitting, we have identified an optimal ”Golden Overfitting Interval” $ARQ_{optimal}$ that maximizes detection efficacy without compromising model reliability.
-This strategic balance enables robust differentiation between normal and anomalous samples, providing a substantial performance boost.
-We also introduce a novel metric, the Relative Anomaly Distribution Index, called RADI, to differentiate from traditional AUROC-pixel measures.
-Unlike AUROC-pixel, RADI is derived from the cumulative distribution function (CDF) of the normal and anomalous pixel prediction score distributions.
-While RADI provides an indirect reflection of detection performance and maintains a similar trend to AUROC-pixel, it also offers a more versatile framework for theoretical modeling and practical discussion of model capabilities.
-This distinction underscores its value as a complementary tool, adding a new layer of analytical depth to our approach.
+Existing anomaly detection metrics such as AUROC-pixel predominantly evaluate performance at discrete thresholds, failing to systematically quantify how overfitting impacts model sensitivity across continuous distribution shifts.
+To fill this critical gap, we introduce the Aberrance Retention Quotient(ARQ), explicitly measuring the extent of overfitting by quantifying prediction deviations from normal data.
+Leveraging ARQ, we define a ”Golden Overfitting Interval” ($ARQ_{optimal}$), wherein controlled overfitting maximizes anomaly detection sensitivity without compromising model stability.
+Additionally, we propose the Relative Anomaly Distribution Index (RADI), an innovative metric derived directly from cumulative distribution functions of normal and anomalous predictions, capturing subtle changes in their overlap and thus providing a more informative assessment than AUROC alone.
 Moreover, we substantiate the rationale for employing Gaussian noise in pseudo-anomaly detection modules.
 Through calculating the Total Variation Distance (TVD) between actual anomaly data and a Gaussian distribution across the entire dataset, we obtain a fair result.
-This outcome indicates a high level of similarity between true anomalies and a Gaussian distribution, thereby providing a theoretical validation for the common practice of using Gaussian noise as an effective pseudo-anomaly generator in previous works like GLASS [4] and SimpleNet [17].
+This outcome indicates a high level of similarity between true anomalies and a Gaussian distribution, thereby providing a theoretical validation for the common practice of using Gaussian noise as an effective pseudo-anomaly generator in previous works like GLASS (Chen et al., 2024) and SimpleNet (Liu et al., 2023).
 This insight not only strengthens our methodology but also offers a theoretical foundation for broader applications in anomaly detection frameworks.
@@
   • We challenge and transcend the conventional understanding of overfitting, positioning it as a controllable and transformative mechanism capable of unlocking model capabilities beyond traditional boundaries.
-  • We introduce the ARQ to precisely regulate the overfitting degree and the RADI as a complementary metric to AUROC-pixel, providing a more adaptable tool for mathematical modeling and facilitating Dual Control Mechanism with the use of COAD.
-  • Our approach dismantles the demonization of overfitting, repurposing it as a generalizable module and achieving SOTA results in both one-class and multi-class anomaly detection tasks.
+  • We introduce the ARQ to precisely regulate the overfitting degree and the RADI as a complementary metric to AUROC-pixel, providing a more adaptable tool for mathematical modeling and facilitating Dual Control Mechanism.
+  • Our method achieves SOTA results in both one-class and multi-class anomaly detection tasks.
 We also provide a theoretical foundation for utilizing Gaussian noise as a preliminarily pseudo-anomaly generator.
-2 Related Work
-2.1 Overview of Anomaly Detection
+2. Related Work
+2.1. Overview of Anomaly Detection
 Reconstruction-Based Method
 Reconstruction-based methods are founded on the hypothesis that models trained exclusively on normal data can successfully reconstruct normal instances while failing to do so for anomalies, thereby using reconstruction error as a measure of abnormality.
-Notable reconstruction-based approaches include Autoencoders (AE)[29, 33], Generative Adversarial Networks (GANs)[7, 15], and Reverse Distillation (RD)[6, 23].
-Specific methods under these categories, such as DSR[31], RealNet[32], and DRAEM[30], attempt to learn the representation of normal data and generate accurate reconstructions for normal samples but tend to produce significant residuals for anomalous inputs.
+Notable reconstruction-based approaches include Autoencoders (AE)(Zavrtanik, 2021; Zhou, 2022), Generative Adversarial Networks (GANs)(Goodfellow et al., 2014; Liang et al., 2023), and Reverse Distillation (RD)(Deng and Li, 2022; Tien et al., 2023).
+Specific methods under these categories, such as DSR(Zavrtanik et al., 2022), RealNet(Zhang et al., 2024), and DRAEM(Zavrtanik et al., 2021), attempt to learn the representation of normal data and generate accurate reconstructions for normal samples but tend to produce significant residuals for anomalous inputs.
 Embedding-Based Method
@@
 Anomalies are identified based on the distance between their embeddings and those of normal samples.
-Techniques such as Double-MMD RGP[25], RealNet[32], PyramidFlow[13], and CFlowAD[9] utilize this paradigm, some of which employ a memory bank to store representative features and apply distance metrics to detect anomalies.
+Techniques such as Double-MMD RGP(Xiao et al., 2023), RealNet(Zhang et al., 2024), PyramidFlow(Lei et al., 2023), and CFlowAD(Gudovskiy et al., 2021) utilize this paradigm, some of which employ a memory bank to store features and apply distance metrics to detect anomalies.
 Embedding-based methods are particularly valuable in high-dimensional feature spaces, where the intricate relationships between different regions of the data can be better captured using a pre-trained model’s learned representations.
@@
 These synthetic anomalies are used to provide explicit anomaly labels in the training set, converting the unsupervised task into a form of supervised learning.
-Notable techniques include CutPaste [14], which pastes cut-out normal patches in different positions, and methods such as GLASS[4], UniAD[28], SimpleNet[17], RealNet[32],and DDPM [12], all of which employ Gaussian noise to generate synthetic anomalies at both feature and image levels.
+Notable techniques include CutPaste (Li et al., 2021), which pastes cut-out normal patches in different positions, and methods such as GLASS(Chen et al., 2024), UniAD(You et al., 2022), SimpleNet(Liu et al., 2023), RealNet(Zhang et al., 2024),and DDPM (Ho et al., 2020), all of which employ Gaussian noise to generate synthetic anomalies at both feature and image levels.
 These approaches help in training the model to differentiate between normal and abnormal samples, thus improving its ability to detect true anomalies.
-By contrast, methods such as DRAEM[30] and DMDD [16] employ Berlin noise or other forms of noise, which have not demonstrated the same level of effectiveness in anomaly detection tasks.
+By contrast, methods such as DRAEM(Zavrtanik et al., 2021) and DMDD (Liu et al., 2024) employ Berlin noise or other forms of noise, which have not demonstrated the same level of effectiveness in anomaly detection tasks.
 In subsequent sections, we will theoretically justify the use of Gaussian noise in these pseudo-anomaly generation modules, establishing its effectiveness and reasonableness as an anomaly synthesis strategy.
-2.2 One-class vs. Multi-class
+2.2. One-class vs. Multi-class
 Our method builds on existing frameworks in anomaly detection and requires a deep understanding of these foundational models.
@@
 One-class anomaly detection focuses on modeling the distribution of normal data to identify deviations.
-Methods such as RD [6] and RD++ [23] are foundational in this area.
+Methods such as RD (Deng and Li, 2022) and RD++ (Tien et al., 2023) are foundational in this area.
 RD employs a pretrain model for feature extraction, while RD++ further enhances this process for improved robustness.
 Multi-class Anomaly Detection
-Multi-class anomaly detection handles diverse object categories and intra-class variations. Notable approaches like UniAD [28] and HVQ-Trans [18] effectively generalize across multiple classes.
-2.3 Non-Diffusion vs. Diffusion-Based
-For the purpose of making our method easier for readers to understand and apply, we have also supplemented it with the diffusion-based framework.
+Multi-class anomaly detection handles diverse object categories and intra-class variations.
+Notable approaches like UniAD (You et al., 2022) and HVQ-Trans (Lu et al., 2023) effectively generalize across multiple classes.
+2.3. Non-Diffusion vs. Diffusion-Based
+For the purpose of making our method easier for readers to apply, we have supplemented it with the diffusion-based framework.
 Non-diffusion methods
-Non-diffusion methods, such as UniAD [28], have been effective for anomaly detection but often face limitations in preserving the semantic consistency of anomalous regions, especially in complex settings.
+Non-diffusion methods, such as UniAD (You et al., 2022) and GLAD (Yao et al., 2024), have been effective for anomaly detection but often face limitations in preserving the semantic consistency of anomalous regions, especially in complex settings.
 Diffusion-based methods
-Diffusion-based methods have recently gained traction due to their exceptional reconstruction capabilities, as evidenced by DiAD [11], which stands out among contemporary methods.
-DiAD employs a diffusion framework for generating anomalies while maintaining semantic information.
-3 Method
-3.1 Overview
-(a) Training Stage
-(b) Inference Stage
-Figure 2: Comprehensive Framework for COAD.(a) Training Stage: In the traditional framework, the teacher model remains frozen throughout.
-The pseudo-anomaly generation and filtering components are optional, remaining unfrozen during the normal training stage but frozen during the controllable overfitting stage, under the guidance of the Dual Control Mechanism.
+Diffusion-based methods have recently gained traction due to their exceptional reconstruction capabilities, as evidenced by DiAD (He et al., 2023), which stands out among contemporary methods.
+Besides, AnomalyDiffusion (Hu et al., 2024) and TransFusion (Fučka et al., 2024), which jointly learn to generate anomalies and discriminate them from normal data by use of Diffusion-based methods.
+3. Method
+3.1. Overview
+Figure 2.
+Comprehensive Framework for COAD.
+In the Traditional Model, Teacher Model remains frozen throughout.
+The pseudo-anomaly Gaussian Noise is optional.
 The student model also remains unfrozen during normal training stage, with selective freezing employed during the controllable overfitting stages.
-When the Dual Control Mechanism fails, Freeze Command Signals are triggered, prompting the selective freezing of certain network layers in the traditional framework to effectively manage overfitting.
-(b) Inference Stage: The inference pipeline follows a similar procedure to the traditional framework, ensuring that the enhancements provided by COAD integrate seamlessly.
+When the Dual Control Mechanism fails, Freeze Command Signals are triggered, prompting the selective freezing of certain network layers and flagging the selected layer in the Traditional Model to effectively manage overfitting.
+The inference pipeline follows a similar procedure to the Traditional Model, ensuring that the enhancements provided by COAD integrate seamlessly.
 As  fig.
-2 shows, our works provides an overview of our proposed framework for enhancing anomaly detection through controllable overfitting.
+2 shows, our work provides an overview of our proposed framework for enhancing anomaly detection through controllable overfitting.
 We introduce Controllable Overfitting-based Anomaly Detection, called COAD.
@@
 These components collectively facilitate the monitoring, control, and optimization of the overfitting process to maximize model sensitivity while preserving generalization capabilities.
-3.2 Quantifying Overfitting with Aberrance Retention Quotient
-The exploration of overfitting has largely been confined to its avoidance, and there has been a conspicuous lack of systematic studies on how to control or even beneficially leverage overfitting for enhancing model capabilities.
-Moreover, existing research lacks a quantitative metric that accurately reflects the extent of overfitting, which is critical for understanding its implications and finding the ”sweet spot” in model training.
-To fill this gap, we introduce the Aberrance Retention Quotient , called ARQ, a novel metric designed to quantify the degree of overfitting by capturing the model’s divergence from true data representation during training.
-The ARQ is formally defined as follows:
-$ARQ=\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})}{\sum_{i=1}^{N}y_{i}},$ (1)
+The detailed overall training and inference pipeline’s pseudo code can be seen in  algorithm 1 and  algorithm 3 in Supplementary Materials.
+Given an input image $x\in\mathcal{X}$, our Teacher network $\mathcal{T}$, which is trained on a large-scale dataset, produces reference features: $y\;=\;\mathcal{T}(x)$.
+The Student network $\mathcal{S}$ then predicts $\hat{y}\;=\;\mathcal{S}(x)$.
+Finally, the discrepancy between $y$ and $\hat{y}$ determines the anomaly region.
+We get the anomaly region $M\;=\;y\;\oplus\;\hat{y}$.
+[…]
+3.2. Quantifying Overfitting with Aberrance Retention Quotient
+The exploration of overfitting has largely been confined to its avoidance, yet systematic methods to quantify, control, and beneficially leverage overfitting remain largely unexplored.
+To fill this gap, we introduce the Aberrance Retention Quotient, called ARQ, a novel metric designed to quantify the degree of overfitting by capturing the model’s divergence from true data representation.
+Specifically, the ARQ is formally defined as follows:
+(1) $ARQ=\frac{\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|}{\sum_{i=1}^{N}y_{i}},$
 where N represents the total number of instances, $\hat{y}_{i}$ represents the predicted output at the $i^{th}$ instance, and $y_{i}$ denotes the corresponding original ground truth value.
@@
 ARQ is a core metric in our framework that quantifies the progression of overfitting during model training.
-By tracking ARQ, we identify the golden overfitting interval, where the model’s sensitivity to anomalous patterns is maximized without compromising generalization.
+Through continuous monitoring of ARQ, we establish an optimal interval golden overfitting interval, where the model’s sensitivity to anomalous patterns is maximized without compromising generalization.
+To objectively determine this interval, we employ a validation-based adaptive search strategy, optimizing ARQ to improve anomaly detection.
 Specifically, we denote the interval of optimal ARQ as:
-$ARQ_{\text{optimal}}=[\theta-\delta,\theta+\delta],$ (2)
+(2) $ARQ_{\text{optimal}}=[\theta-\delta,\theta+\delta],$
 where $\theta$ is the baseline value around which the optimal range is established, $\delta$ is a quantity of the same order of magnitude as $\theta$, representing the permissible deviation from the baseline $\theta$ within the optimal range, which indicates the region in which the overfitting is leveraged most effectively.
@@
 By effectively using ARQ, we transform overfitting into a controlled mechanism that enhances model robustness and discriminative power for anomaly detection.
-3.3 Bridging Theoretical Modeling and Practical Anomaly Detection with Relative Anomaly Distribution Index
-3.3.1 Probability Distribution Models of Normal and Anomalous Pixel Scores
-To facilitate the mathematical treatment later, we made the following two assumptions, which were verified in subsequent experiments, please see  Sec. 4.
+3.3. Bridging Theoretical Modeling and Practical Anomaly Detection with Relative Anomaly Distribution Index
+3.3.1. Probability Distribution Models of Normal and Anomalous Pixel Scores
+To facilitate the mathematical treatment later, we make the following two assumptions.
+We empirically verify these assumptions and find that normal distributions provide adequate approximations for prediction score distributions, a common phenomenon justified by central limit theorem and supported by empirical observations (refer to  fig.
+5 and Supplementary Material).
 Distribution of Normal Pixel Prediction Scores
 At an Aberrance Retention Quotient of $\textit{ARQ}=\theta$, the prediction scores of normal pixels $S_{n}$ follow a normal distribution dependent on $\theta$:
-$S_{n}\sim N(\mu_{n}(\theta),\sigma_{n}(\theta)^{2}),$ (3)
+(3) $S_{n}\sim N(\mu_{n}(\theta),\sigma_{n}(\theta)^{2}),$
 where the mean $\mu_{n}(\theta)$ and variance $\sigma_{n}(\theta)^{2}$ vary with ARQ. As the ARQ increases, the model’s memory of normal samples is enhanced, potentially resulting in:
@@
 The variance of normal pixel prediction scores decreases exponentially with increasing ARQ, expressed as:
-$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta},$ (4)
+(4) $\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta},$
 where $\sigma_{n0}$ is the initial variance (without overfitting), and $k$ is a positive constant representing the rate at which variance decreases.
@@
 The prediction scores of anomalous pixels $S_{a}$ follow a fixed normal distribution that does not change with ARQ:
-$S_{a}\sim N(\mu_{a},\sigma_{a}^{2}),$ (5)
-since anomalous pixels are not present during training, and the model cannot memorize them.
-3.3.2 Introduction of RADI
+(5) $S_{a}\sim N(\mu_{a},\sigma_{a}^{2}),$
+Since anomalous pixels represent unseen deviations that the model does not explicitly memorize during training, we assume their predicted score distribution remains relatively stable and unaffected by varying degrees of overfitting(refer to  fig.
+5).
+3.3.2. Introduction of RADI
 After these assumptions, we introduce the Relative Anomaly Distribution Index, called RADI, a novel metric designed to bridge theoretical modeling and practical anomaly detection.
-Unlike traditional metrics such as AUROC, which are often used to evaluate model performance in anomaly detection tasks, RADI provides an indirect yet insightful measure of the model’s capacity to differentiate between normal and anomalous samples.
-This makes it particularly valuable for theoretical analyses, while also retaining practical utility in assessing model effectiveness.
-The RADI quantifies the overlap between normal and anomalous score distributions, akin to the Wilcoxon rank-sum [24] and Mann-Whitney U tests [19].
+Unlike threshold-dependent metrics such as AUROC, RADI directly measures the probabilistic separation between normal and anomalous distributions, providing more explicit insights into the model’s discriminative capability independent of specific thresholds.
+This makes it particularly valuable for theoretical analyses, while retaining practical utility in assessing model effectiveness.
+The RADI quantifies the overlap between normal and anomalous score distributions, akin to the Wilcoxon rank-sum (Wilcoxon, 1992) and Mann-Whitney U tests (Mann and Whitney, 1947).
 RADI is calculated using the cumulative distribution function (CDF) of the score distributions:
@@
 where $S_{a}$ and $S_{n}$ represent the scores of anomalous and normal pixels, respectively.
-It might seem counterintuitive why using RADI as an indicator would enhance model performance, as one might think that maximizing this value would lead to the model predicting everything as anomalous.
-However, RADI itself is not directly used for model optimization; rather, it serves as a metric to reflect the optimization outcome and help identify the extent to which normal and anomalous distributions are distinguishable.
-The goal is to ensure that anomalies receive consistently higher scores compared to normal samples, which allows the model to achieve better discrimination.
-Moreover, the relationship between overfitting and variance is controlled by  Eq.
-8, which ensures that the variance of normal pixel predictions ($\sigma_{n}(\theta)$) does not decrease to zero due to the noise term $\sigma_{\text{noise}}(\theta)$.
-This prevents the variance from collapsing entirely, maintaining a distinction between normal and anomalous scores.
-Thus, RADI can increase without the model predicting all samples as anomalous, as variability in normal predictions remains above a non-zero threshold, sustaining a balance between sensitivity to anomalies and generalization.
+Intuitively, RADI quantifies the probability that a selected anomalous pixel has a higher anomaly score than a selected normal pixel.
+A higher RADI thus indicates clearer model discrimination between anomalies and normal patterns.
+Although maximizing RADI might seem to encourage over-classifying anomalies, in practice, the variance of normal pixel scores remains above a non-zero lower bound due to controlled noise ( eq.
+8).
+This ensures normal samples retain distinct scores from anomalies, preventing the trivial solution of classifying all samples as anomalous.
 RADI provides a probabilistic assessment of whether anomalous scores exceed normal scores, offering insights into model discrimination.
@@
 Overall, RADI complements AUROC by providing a simplified measure of the model’s discriminative power, effectively reflecting its trend without requiring threshold-based evaluations.
-3.3.3 Theoretical Derivation of Optimal Overfitting Conditions
-For our case, where both $S_{a}$ and $S_{n}$ follow normal distributions, RADI can be further expressed in a simplified analytical form:
-$\textit{RADI}(\theta)=\Phi\left(\frac{\mu_{a}-\mu_{n}}{\sqrt{\sigma_{n}(\theta)^{2}+\sigma_{a}^{2}}}\right),$ (7)
+3.3.3. Theoretical Derivation of Optimal Overfitting Conditions
+For our case, where both $S_{a}$ and $S_{n}$ follow normal distributions, to analytically derive the optimal level of controllable overfitting, we explicitly express RADI in closed-form under normal distribution assumptions.
+This facilitates understanding how statistical changes induced by overfitting influence anomaly detection performance:
+(7) $\textit{RADI}(\theta)=\Phi\left(\frac{\mu_{a}-\mu_{n}}{\sqrt{\sigma_{n}(\theta)^{2}+\sigma_{a}^{2}}}\right),$
 where $\Phi$ is the CDF of the standard normal distribution.
-In this case, $\mu_{a}$ and $\mu_{n}$ are the means of the anomaly and normal pixel score distributions, respectively, while $\sigma_{a}$ and $\sigma_{n}(\theta)$ are their respective standard deviations, with $\theta$ representing the Aberrance Retention Quotient.
+In this case, $\mu_{a}$ and $\mu_{n}$ are the means of the anomaly and normal pixel score distributions, respectively, while $\sigma_{a}$ and $\sigma_{n}(\theta)$ are their respective standard deviations, with $\theta$ representing the ARQ.
 eq.
@@
 To describe this phenomenon, we introduce a noise term:
-$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta}+\sigma_{\text{noise}}(\theta),$ (8)
+(8) $\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta}+\sigma_{\text{noise}}(\theta),$
 where $\sigma_{\text{noise}}(\theta)$ represents the noise impact induced by overfitting and can be expressed as:
-$\sigma_{\text{noise}}(\theta)=\sigma_{\text{max}}(1-e^{-h(\theta-\theta_{0})}),\quad\text{for}\;\theta>\theta_{0},$ (9)
+(9) $\sigma_{\text{noise}}(\theta)=\sigma_{\text{max}}(1-e^{-h(\theta-\theta_{0})}),\quad\text{for}\;\theta>\theta_{0},$
 where $\sigma_{\text{max}}$ is the maximum noise variance, $h$ controls the rate of increase of noise variance, and $\theta_{0}$ is the threshold of ARQ where noise starts to appear.
 When $\theta\leq\theta_{0}$, $\sigma_{\text{noise}}(\theta)=0$.
-Category DSR [31] PatchCore [21] BGAD [26] SimpleNet [17] GLASS [4] RD [6] RD (+COAD) RD++ [23] RD++ (+COAD)
-Object Bottle 99.6/98.8 100/98.5 100/98.9 100/98.0 100/99.3 100/98.7 100/98.8 100/98.8 100/99.6
+Table 1.
+Performance comparison across different methods in one-class tasks on MVTec AD dataset.
+Bold text indicates the best performance among all methods, while underlined text shows the comparison Method vs.
+Method (+COAD).
+The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
+Category DSR (Zavrtanik et al., 2022) PatchCore (Roth et al., 2022) BGAD (Yao et al., 2023) SimpleNet (Liu et al., 2023) GLASS (Chen et al., 2024) RD (Deng and Li, 2022) RD (+COAD) RD++ (Tien et al., 2023) RD++ (+COAD)
+Object
+Bottle 99.6/98.8 100/98.5 100/98.9 100/98.0 100/99.3 100/98.7 100/98.8 100/98.8 100/99.6
 Cable 95.3/97.7 99.8/98.4 97.9/98.0 100/97.5 99.8/98.7 95.0/97.3 97.9/97.7 99.3/98.4 99.4/99.2
@@
 Zipper 98.5/98.4 99.4/98.9 97.8/98.7 99.9/98.9 100/99.6 98.5/98.2 99.1/98.7 98.6/98.8 99.8/99.5
-Texture Carpet 99.6/96.0 98.6/99.1 99.8/99.4 99.7/98.4 99.8/99.6 98.9/98.9 99.8/99.2 100/99.2 99.8/99.5
+Texture
+Carpet 99.6/96.0 98.6/99.1 99.8/99.4 99.7/98.4 99.8/99.6 98.9/98.9 99.8/99.2 100/99.2 99.8/99.5
 Grid 100/99.6 97.7/98.8 99.1/99.4 99.9/98.5 100/99.4 100/99.1 100/99.3 100/99.3 100/99.5
@@
 Avg. 98.2/95.8 99.1/98.1 97.9/98.2 99.5/98.1 99.9/99.3 98.5/97.7 99.4/98.0 99.4/98.3 99.9/99.4
-Table 1: Performance comparison across different methods in one-class tasks.
-Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method.
-The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
+Table 2. Performance comparison across different methods in one-class tasks on VisA dataset.
+Category PatchCore (Roth et al., 2022) SimpleNet (Liu et al., 2023) DRAEM (Zavrtanik et al., 2021) RD (Deng and Li, 2022) RD(+COAD) RD++ (Tien et al., 2023) RD++(+COAD)
+Single Instance Cashew 97.7/99.2 94.8/99.0 88.3/85.0 99.6/95.8 97.8/97.6 97.8/95.8 98.1/98.3
+Chewinggum 99.1/98.9 100/98.5 96.4/97.7 99.7/99.0 97.8/99.4 96.4/99.4 97.4/99.5
+Fryum 91.6/95.9 96.6/94.5 94.7/82.5 96.6/94.3 95.5/96.3 95.8/96.5 96.8/98.1
+Pipe fryum 99.0/99.3 99.2/99.3 94.7/65.8 97.0/92.0 99.4/99.1 99.6/99.1 99.3/99.4
+Multiple Instances Candle 98.7/99.2 96.9/98.6 89.6/91.0 92.2/97.9 95.1/99.0 96.4/98.6 92.8/98.7
+Capsules 68.8/96.5 89.5/99.2 89.2/99.0 90.1/89.5 87.5/99.3 92.1/99.4 92.7/99.7
+Macaroni1 90.1/98.5 97.6/99.6 93.9/99.4 98.4/97.7 95.1/99.3 94.0/99.7 95.7/99.9
+Macaroni2 63.4/93.5 83.4/96.4 88.3/99.7 97.6/87.7 96.6/99.3 88.0/99.7 92.5/99.6
+Complex Structure PCB1 96.0/99.8 99.2/99.8 84.7/98.4 97.6/75.0 96.5/99.7 97.0/99.7 98.8/99.5
+PCB2 95.1/98.4 99.2/98.8 96.2/94.0 91.1/64.8 94.0/98.8 97.2/99.0 97.3/98.9
+PCB3 93.0/98.9 98.6/99.2 97.4/94.3 95.5/95.5 95.9/99.3 96.8/99.2 97.4/99.3
+PCB4 99.5/98.3 98.9/98.6 98.8/97.6 96.5/92.8 99.9/98.7 99.8/98.6 98.5/99.6
+Avg. 91.0/98.1 96.2/98.5 92.4/92.0 96.0/90.1 95.9/98.8 95.9/98.7 96.4/99.2
 To find the optimal Aberrance Retention Quotient of $\textit{ARQ}=\theta^{*}$ that maximizes RADI, we calculate the derivative of $\text{AUROC}_{\text{pixel}}(\theta)$ and set it to zero.
 The derivative is as below and we set the derivative to zero:
-$\frac{d\textit{RADI}(\theta)}{d\theta}=\phi(z(\theta))\cdot\frac{dz(\theta)}{d\theta}=0,$ (10)
+(10) $\frac{d\textit{RADI}(\theta)}{d\theta}=\phi(z(\theta))\cdot\frac{dz(\theta)}{d\theta}=0,$
 where $\phi(z)$ is the probability density function of the standard normal distribution.
-$\frac{dz(\theta)}{d\theta}=-\frac{(\mu_{a}-\mu_{n})\cdot\sigma_{n}(\theta)\cdot\frac{d\sigma_{n}(\theta)}{d\theta}}{(\sigma_{n}(\theta)^{2}+\sigma_{a}^{2})^{3/2}}=0,$ (11)
+(11) $\frac{dz(\theta)}{d\theta}=-\frac{(\mu_{a}-\mu_{n})\cdot\sigma_{n}(\theta)\cdot\frac{d\sigma_{n}(\theta)}{d\theta}}{(\sigma_{n}(\theta)^{2}+\sigma_{a}^{2})^{3/2}}=0,$
 This implies either $\sigma_{n}(\theta)=0$ (unrealistic as variance cannot be zero) or $\frac{d\sigma_{n}(\theta)}{d\theta}=0$. Thus, solving yields:
-$\theta^{*}=\frac{\ln(k\sigma_{n0})-\ln(h\sigma_{\text{max}})+h\theta_{0}}{k-h}.$ (12)
-3.3.4 Dual Control Mechanism and Freeze Command Signals
+(12) $\theta^{*}=\frac{\ln(k\sigma_{n0})-\ln(h\sigma_{\text{max}})+h\theta_{0}}{k-h}.$
+3.3.4. Dual Control Mechanism and Freeze Command Signals
 We define Aberrance Retention Quotient of $\textit{ARQ}=\theta$, adjusted to optimize the pixel-level AUROC.
 The mathematical relationship between ARQ and AUROC is established by monitoring the gradient, we call it Gradient-Guided Overfitting Control.
-$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0.$ (13)
+The Dual Control Mechanism leverages both ARQ and RADI simultaneously: ARQ ensures overfitting is maintained within beneficial bounds, while RADI’s gradient condition ensures the continual improvement of anomaly detection performance.
+(13) $\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0.$
 We integrate $ARQ_{optimal}$ in  eq. 2 and Gradient-Guided Overfitting Control in  eq. 13 and call this combined control mechanism the Dual Control Mechanism:
-$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0,\quad\text{for}\;\theta\in\textit{$ARQ_{optimal}$}.$ (14)
-If Dual Control Mechanism fails, the Dual Control Mechanism triggers Freeze Command Signals, which prompt whether freeze certain network layers, beginning with lower-level feature extraction layers and moving upwards.
+(14) $\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0,\quad\text{for}\;\theta\in\textit{$ARQ_{optimal}$}.$
+When ARQ exceeds its optimal bounds and RADI gradient condition fails, Freeze Command Signals are activated, progressively freezing layers starting from the earliest layers upwards.
 The rationale for progressively freezing certain layers while overfitting the other parts lies in the architecture’s functional partitioning.
@@
 Conversely, keeping deeper layers unfrozen during the overfitting stage enables the model to specialize in identifying unique, high-level features that are more discriminative between normal and anomalous samples.
-This approach helps maintain generalization while enhancing sensitivity to rare patterns, ultimately boosting anomaly detection performance without succumbing to the pitfalls of overfitting the entire model.
-4 Experiments
-Category US [2] PSVDD [27] PaDiM [5] MKD [22] DRAEM [30] UniAD [28] UniAD (+COAD) DiAD [11] DiAD (+COAD)
-Object Bottle 84.0/67.9 85.5/86.7 97.9/96.1 98.7/91.8 97.5/87.6 99.7/98.1 100/98.0 99.7/98.4 100/98.8
+This approach ultimately boosts performance without succumbing to the pitfalls of overfitting the entire model.
+4. Experiments
+Table 3. Performance comparison across different methods in multi-class tasks on MVTec AD dataset.
+Category US (Bergmann et al., 2020) PSVDD (Yi and Yoon, 2020) PaDiM (Defard et al., 2021) MKD (Salehi et al., 2021) DRAEM (Zavrtanik et al., 2021) UniAD (You et al., 2022) UniAD (+COAD) DiAD (He et al., 2023) DiAD (+COAD)
+Object
+Bottle 84.0/67.9 85.5/86.7 97.9/96.1 98.7/91.8 97.5/87.6 99.7/98.1 100/98.0 99.7/98.4 100/98.8
 Cable 60.0/78.3 64.4/62.2 70.9/81.0 78.2/89.3 57.8/71.3 95.2/97.3 97.9/98.8 94.8/96.8 97.5/97.7
@@
 Zipper 78.6/84.2 77.8/95.1 79.7/94.8 87.4/86.1 98.8/98.3 95.8/96.8 97.6/98.1 95.1/96.2 98.3/98.7
-Texture Carpet 86.6/88.7 63.3/78.6 93.8/97.6 69.8/95.5 98.0/98.6 99.8/98.5 100/98.8 99.4/98.6 99.6/99.2
+Texture
+Carpet 86.6/88.7 63.3/78.6 93.8/97.6 69.8/95.5 98.0/98.6 99.8/98.5 100/98.8 99.4/98.6 99.6/99.2
 Grid 69.2/64.5 66.0/70.8 73.9/71.0 83.8/82.3 99.3/98.7 98.2/96.5 99.1/98.4 98.5/96.6 100/99.3
@@
 Avg. 74.5/81.8 76.8/85.6 84.2/89.5 81.9/84.9 88.1/87.2 96.5/96.8 97.6/98.0 97.2/96.8 99.1/98.0
-Table 2: Performance comparison across different methods in multi-class tasks.
-Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method.
-The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
-4.1 Dataset
-We evaluate our method using the MVTec AD dataset [1], which is a comprehensive, real-world dataset for unsupervised anomaly detection tasks.
+We evaluate our method using the MVTec AD dataset (Bergmann et al., 2019) and VisA dataset (Zou et al., 2022), which are comprehensive, real-world datasets for unsupervised anomaly detection tasks.
 The MVTec AD dataset includes 15 categories, comprising 3629 images for training and 1725 images for testing.
-The dataset provides pixel-level annotations for each anomaly, making it suitable for evaluating both image-level classification and pixel-level segmentation tasks.
-4.2 Comparison with SOTAs
-In this section, we compare the performance of our proposed method with several SOTA anomaly detection methods in both one-class and multi-class tasks.
+The VisA dataset includes 12 categories, including 9,621 normal and 1,200 anomalous samples.
+4.1. Comparison with SOTAs
+In this section, we compare the performance of our proposed method with several SOTA anomaly detection methods in both one-class and multi-class tasks on MVTec AD and VisA datasets.
 Using the optimal ARQ value $\theta$ derived from the $ARQ_{optimal}$ as indicated by  eq.
-12, we set ARQ to 0.006 for one-class tasks and 0.06 for multi-class tasks.
+12, we set ARQ to 0.006 for one-class tasks and 0.06 for multi-class tasks as shown in  section 4.2.2 which is chosen empirically.
 $ARQ_{optimal}$ is defined as $\theta\in[0.001,0.011]$ for one-class tasks and $\theta\in[0.01,0.11]$ for multi-class tasks, where the balance between overfitting and generalization is optimal for anomaly detection.
-We saved model states within these intervals and evaluated their performance across respective benchmarks to ensure robust detection.
-Furthermore, during model training, we adopted a Dual Control Mechanism for Optimal Detection Performance approach in our framework COAD, as detailed in  Eq.
-14, and when Dual Control Mechanism fails, it triggers Freeze Command Signals, prompting the selective freezing of certain network layers in the traditional framework.
+Furthermore, during model training, we adopt a Dual Control Mechanism for Optimal Performance method in our framework COAD, as detailed in  eq.
+14, and when Dual Control Mechanism fails, it triggers Freeze Command Signals, prompting selective freezing of certain network layers in the traditional framework.
 Our method builds upon existing anomaly detection frameworks.
@@
 After training, the inference of the model proceeds in the same manner as the baseline anomaly detection frameworks, ensuring that the enhancements provided by COAD integrate seamlessly.
-The results of our one-class and multi-class experiments are summarized in  Tabs.
-1 and 2.
-It can be observed that our enhanced method outperforms the original methods, achieving new SOTA results.
-The enhancement in pixel-level AUROC is especially significant, which is crucial for fine-grained anomaly detection.
-These results confirm that the introduction of controllable overfitting, regulated by ARQ and RADI, enhances the model’s sensitivity and stability, leading to robust anomaly detection abilities that outperform existing SOTA methods.
-4.3 Ablation Study
-4.3.1 Comparisons with Our Basic Frameworks
+The quantitative results of our experiments across both one-class and multi-class anomaly detection tasks are comprehensively summarized in  tables 1, 2 and 3, with additional detailed results for the VisA multi-class task provided in  table 9 of the Supplementary Materials  appendix D.
+Clearly, our proposed COAD framework consistently surpasses previous SOTA approaches, delivering substantial improvements in both image-level and pixel-level AUROC metrics across all evaluated datasets and tasks.
+Additionally, partial qualitative visualization results demonstrating the effectiveness of COAD are presented in  figs.
+3 and 4, showing high accuracy.
+Complete qualitative comparisons are provided in  appendix E of the Supplementary Materials.
+Figure 3. Qualitative visualization results on partial categories from the MVTec AD for the one-class task. Columns show Original images, GroundTruth, and predictions generated by our COAD framework.
+Figure 4. Qualitative visualization results on partial categories from the MVTec AD for the multi-class task.
+In conclusion, these comprehensive experimental validations solidify the practical advantages of our COAD framework, highlighting its ability to reliably enhance existing anomaly detection models through controlled and strategic utilization of overfitting.
+4.2. Ablation Study
+4.2.1. Comparisons with Our Basic Frameworks
+Table 4. Image-Level AUROC comparison between the basic and enhanced versions of different frameworks.
 RD RD++ UniAD DiAD
-Baseline 98.5 99.4 96.5 97.2
+MVTec AD Baseline 98.5 99.4 96.5 97.2
 +COAD 99.4(+0.9) 99.9(+0.5) 97.6(+1.1) 99.1(+1.9)
-Table 3: Image-Level AUROC comparison between the basic and enhanced versions of different frameworks.
+VisA Baseline 96.0 95.9 85.5 -
++COAD 95.9(-0.1) 96.4(+0.5) 91.0(+5.5) -
+Table 5. Pixel-Level AUROC comparison between the basic and enhanced versions of different frameworks.
 RD RD++ UniAD DiAD
-Baseline 97.7 98.3 96.8 96.8
+MVTec AD Baseline 97.7 98.3 96.8 96.8
 +COAD 98.0(+0.3) 99.4(+1.1) 98.0(+1.2) 98.0(+1.2)
-Table 4: Pixel-Level AUROC comparison between the basic and enhanced versions of different frameworks.
-We conduct an ablation study comparing our enhanced framework with its corresponding basic versions in  Tabs.
-3 and 4.
+VisA Baseline 90.1 98.7 95.9 -
++COAD 98.8(+8.7) 99.2(+0.5) 97.5(+1.6) -
+We conduct an ablation study comparing our enhanced framework with its corresponding basic versions in  tables 4 and 5.
 The results demonstrate consistent and notable improvements in both image-level and pixel-level AUROC across all evaluated frameworks.
 In particular, the performance gains observed in image-level anomaly detection emphasize the strength of our COAD, which effectively captures subtle differences between normal and anomalous regions.
-Specifically, our enhancements yielded an increase of up to 1.9 in the DiAD framework, as seen in image-level AUROC, and similar significant improvements in other frameworks.
-For pixel-level AUROC, a similar pattern of improvement was observed, with gains of up to 1.2 seen in multiple frameworks, as shown in Tab.
-4.
+Specifically, our enhancements yield an increase of up to 1.9 in the DiAD framework, as seen in image-level AUROC, and similar significant improvements in other frameworks.
+For pixel-level AUROC, a similar pattern of improvement is observed, with gains of up to 1.2 seen in multiple frameworks, as shown in table 5.
 This underscores the strength of our COAD, which not only optimizes image-level detection but also provides enhanced granularity in distinguishing pixel-level abnormalities.
 Furthermore, the versatility of our method is evident in its applicability to different anomaly detection paradigms.
-Whether in one-class or multi-class anomaly detection tasks, or even in settings employing diffusion-based models versus non-diffusion frameworks, our method has consistently demonstrated superior performance, showcasing our method’s robustness and generalizability.
-4.3.2 Comparisons with Distinct ARQ
+Our method has consistently demonstrated superior performance, showcasing our method’s robustness and generalizability.
+4.2.2. Comparisons with Distinct ARQ
+Table 6. Comparison of performance across distinct ARQ ranges for different models in one-class tasks on MVTec AD.
 Model 0–0.001 0.001–0.011 0.011–0.1
@@
 RD++ 99.7 / 98.8 99.9 / 99.4 -
-Table 5: Comparison of performance across distinct ARQ ranges for different models in one-class tasks.
+Table 7. Comparison of performance across distinct ARQ ranges for different models in multi-class tasks on MVTec AD.
 Model 0–0.01 0.01–0.11 0.11–1
@@
 DiAD 97.3 / 96.6 99.1 / 98.0 99.2 / 96.5
-Table 6: Comparison of performance across distinct ARQ ranges for different models in multi-class tasks.
 In this section, we compare the performance of different models under distinct ARQ ranges to validate the impact of controllable overfitting on anomaly detection.
-We evaluated methods such as RD, RD++, UniAD, and DiAD across various ARQ levels.
-As  Tabs.
-5 and 6 show, the results demonstrate that, by appropriately regulating the ARQ, which is in $ARQ_{optimal}$, model detection performance can be significantly enhanced.
-It is evident that our controllable overfitting strategy not only yields consistent improvements across different models but also helps in striking the optimal balance between overfitting and generalization.
+We evaluate methods such as RD, RD++, UniAD, and DiAD across various ARQ levels.
+As  tables 6 and 7 show, the results demonstrate that, by appropriately regulating the ARQ, which is in $ARQ_{optimal}$, model detection can be significantly enhanced.
+It is evident that our controllable overfitting strategy not only yields consistent improvements across different models but also helps in striking the balance between overfitting and generalization.
 Overall, these results demonstrate that our controllable overfitting strategy consistently yields significant performance gains across various models and tasks.
 By leveraging the ARQ within the identified optimal range, we successfully enhance the models’ discriminative abilities while maintaining a balanced relationship between overfitting and generalization.
-4.4 Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
+4.3. Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
+(a) Carpet (Normal)
+(b) Wood (Normal)
+(c) Carpet (Anomaly)
+(d) Wood (Anomaly)
+Figure 5. Histograms for normal and anomaly pixel scores of partial categories of MVTec AD dataset. Each histogram includes Gaussian fitting curves for better visualization.
+Table 8. Partial Results of Validation Metrics for Assumptions on Normal and Anomalous Pixel Score Distributions.
 ARQ $\mu_{n}$ $\theta_{n}$ $\mu_{a}$ $\theta_{a}$ $TVD_{n}$ $TVD_{a}$
@@
 0.082 105.5 65.0 99.0 45.0 0.19 0.08
-Table 7: Partial Results of Validation Metrics for Assumptions on Normal and Anomalous Pixel Score Distributions.
-In this section, we firstly validate the assumptions introduced in Sec.
-3.3.1 by employing the Total Variation Distance (TVD) metric.
-We leveraging multiple models at different stages of overfitting, characterized by distinct ARQ values.
-For each model state, corresponding to different ARQ = $\theta$, we calculate the prediction scores for both normal and anomalous pixels across the dataset.
-Subsequently, we evaluate their proximity to a Gaussian distribution using TVD, ensuring a robust assessment of the normality of these distributions.
-The partial results of this validation are summarized in Tab.
-7.
-Moreover, we model the variance of the normal pixel prediction scores $(\sigma_{n}(\theta))$ as an exponential function of the ARQ = $\theta$, given by Eq.
+In this section, we firstly validate the assumptions introduced in section 3.3.1 by employing the Total Variation Distance (TVD) metric.
+We leverage multiple models at different stages of overfitting, characterized by distinct ARQ values.
+For each model state, corresponding to different ARQ = $\theta$, we calculate the prediction scores for all pixels across the dataset.
+Subsequently, we evaluate their proximity to a Gaussian distribution using TVD.
+The partial results of this validation are summarized in table 8.
+Moreover, we model the variance of the normal pixel prediction scores $(\sigma_{n}(\theta))$ as an exponential function of the ARQ, given by eq.
 4.
+Partial qualitative result can be seen in  fig.
+5 and complete qualitative study can be seen in  figs.
+7 and 6 in Supplementary Materials.
 To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we employ the TVD metric to quantitatively compare the distribution of real anomaly scores with that of Gaussian-distributed pseudo-anomalies.
-We can see  Tab.
-7, the resulting TVD value was about 0.08, suggesting that the Gaussian-distributed pseudo-anomalies exhibit a significant overlap with the actual anomalies.
-This supports the assumptions that Gaussian noise can be effectively used to preliminarily simulate anomalous behavior in training models for anomaly detection, which is usually used by GLASS[4], SimpleNet[17], RealNet[32],and DDPM [12].
-5 Conclusion
-In this work, we reenvisioned overfitting as a controllable and transformative mechanism, using COAD to enhance model capabilities beyond conventional boundaries.
-We introduced ARQ to precisely regulate overfitting, as well as RADI, which leverages $ARQ_{optimal}$ to provide a more versatile metric compared to AUROC-pixel, facilitating both theoretical modeling and Dual Control Mechanism.
+We can see  table 8, the resulting TVD value is about 0.08, suggesting that the Gaussian-distributed pseudo-anomalies exhibit a significant overlap with the actual anomalies.
+This supports the assumptions that Gaussian noise can be used to preliminarily simulate anomalous behavior for anomaly detection, which is usually used by GLASS(Chen et al., 2024), SimpleNet(Liu et al., 2023), RealNet(Zhang et al., 2024),and DDPM (Ho et al., 2020).
+Extended Discussion of Gaussian Noise Validation can be seen in  appendix C in Supplementary Materials.
+5. Conclusion
+In this work, we reconceptualize overfitting as a controllable and transformative mechanism, using COAD to enhance model capabilities beyond conventional boundaries.
+We introduce ARQ to precisely regulate overfitting, as well as RADI, which leverages $ARQ_{optimal}$ to provide a more versatile metric compared to AUROC-pixel, facilitating both theoretical modeling and Dual Control Mechanism.
 By repurposing overfitting as a generalizable module, our approach not only dismantles its demonization but also enables us to achieve SOTA results in both one-class and multi-class anomaly detection tasks.
 Furthermore, we provide a robust theoretical foundation for employing Gaussian noise as a preliminary pseudo-anomaly generator, extending the applicability of our COAD.
-Supplementary Material
-6 Dual Control Mechanism Design Concept
+Appendix
+Appendix A Dual Control Mechanism Design Concept
 Overfitting has traditionally been viewed as a problem that hampers generalization—where the model memorizes noise and specific details of the training data, leading to poor performance on unseen examples.
@@
 By monitoring both ARQ and the RADI gradient, and employing Freeze Command Signals when necessary, our framework transforms overfitting from a challenge into an asset—amplifying the model’s ability to detect anomalies while maintaining robustness and generalization.
-7 Main Functional Pseudocode
+Appendix B Main Functional Pseudocode
 Algorithm 1 Training Stage with Gaussian Noise, Teacher-Student Reconstruction, and Overfitting Control
@@
 0: Predicted values $\hat{y}_{i}$, Ground truth values $y_{i}$, Total instances $N$
-1:   Initialize ARQ value: $\textit{ARQ}\leftarrow 0$
+[…]
 2:Compute Numerator:
@@
 0: Scores of anomalous pixels $S_{a}$, Scores of normal pixels $S_{n}$
-1:   Initialize RADI value: $\textit{RADI}\leftarrow 0$
+[…]
 2:Compute RADI using Integral:
@@
 4:Output: Final RADI value
-8 Extended Discussion of Gaussian Noise Validation
-8.1 Theoretical Rationale for Gaussian Noise
-In this work, we chose to use Gaussian noise as a preliminarily pseudo-anomaly generator due to its statistical properties and simplicity.
+Appendix C Extended Discussion of Gaussian Noise Validation
+C.1. Theoretical Rationale for Gaussian Noise
+In this work, we choose to use Gaussian noise as a preliminarily pseudo-anomaly generator due to its statistical properties and simplicity.
 Gaussian noise, characterized by its mean and variance, serves as an ideal candidate for generating random deviations that can mimic unexpected patterns in normal data.
@@
 This universality underpins our approach to controllable overfitting within the COAD framework.
-8.2 Validation Through Total Variation Distance (TVD)
-To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we conducted an empirical analysis using the Total Variation Distance (TVD) metric.
+C.2. Validation Through Total Variation Distance (TVD)
+To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we conduct an empirical analysis using the Total Variation Distance (TVD) metric.
 TVD is a standard measure used in probability theory to quantify the dissimilarity between two distributions.
-We calculated TVD to evaluate how closely Gaussian noise could mimic true anomalies in both spatial and intensity domains.
+We calculate TVD to evaluate how closely Gaussian noise could mimic true anomalies in both spatial and intensity domains.
 The calculation of TVD involves estimating the difference between the empirical distributions of Gaussian noise and the true normal and anomaly pixel intensities:
-$D_{\text{TVD}}(P,Q)=\frac{1}{2}\sum_{x}|P(x)-Q(x)|,$ (15)
+(15) $D_{\text{TVD}}(P,Q)=\frac{1}{2}\sum_{x}|P(x)-Q(x)|,$
 where $P$ represents the empirical distribution of the true anomalies, and $Q$ represents the empirical distribution of Gaussian noise.
 TVD measures the extent of overlap between the two distributions, where a smaller TVD implies a closer approximation between the pseudo-anomalies and real anomalies.
-We conducted experiments on our dataset using both generated Gaussian noise and real anomalies.
+We conduct experiments on our dataset using both generated Gaussian noise and real anomalies.
 presents a comparison of the TVD values for different classes.
 The results indicate that the Gaussian distribution closely matches the distribution of real normal and anomalies across a range of classes.
-Specifically, it can be seen in  Sec.
-4.4.
-8.3 Experimental Results and Qualitative Study
+Specifically, it can be seen in  fig.
+5.
+Table 9. Performance comparison across different methods in multi-class tasks on VisA dataset.
+Category DRAEM (Zavrtanik et al., 2021) JNLD (Liu et al., 2023) OmniAL (Zhao, 2023) UniAD (You et al., 2022) UniAD(+COAD)
+Single Instance Cashew 67.3 / 68.5 82.5 / 94.1 88.6 / 95.0 92.8/98.6 93.7/98.6
+Chewing gum 90.0 / 92.7 96.0 / 98.9 96.4 / 99.0 96.3/98.8 96.6/99.0
+Fryum 86.2 / 83.2 91.9 / 90.0 94.6 / 92.1 83.0/95.9 87.9/97.3
+Pipe fryum 87.1 / 72.3 87.5 / 92.5 86.1 / 98.2 94.7/98.9 97.8/98.4
+Multiple Instances Candle 70.2 / 82.6 85.4 / 94.5 86.8 / 95.8 94.1/98.5 95.8/98.7
+Capsules 89.6 / 96.6 91.4 / 99.6 90.6 / 99.4 55.6/88.7 78.3/97.1
+Macaroni 1 68.6 / 89.8 90.5 / 93.3 92.6 / 98.6 79.9/97.4 87.2/98.2
+Macaroni 2 60.3 / 83.2 71.3 / 92.1 75.2 / 97.9 71.6/95.2 80.4/97.6
+Complex Structure PCB1 83.9 / 94.0 82.9 / 98.0 77.7 / 97.6 92.8/93.3 96.9/97.1
+PCB2 81.7 / 94.1 79.1 / 95.0 81.0 / 93.9 87.8/93.9 94.1/95.2
+PCB3 87.7 / 94.1 90.1 / 98.5 88.1 / 94.7 78.6/97.3 84.5/98.0
+PCB4 87.1 / 72.3 96.2 / 97.5 95.3 / 97.1 98.8/94.9 98.7/94.7
+Avg. 80.0/85.3 87.1/95.3 87.8/96.6 85.5/95.9 91.0/97.5
+C.3. Experimental Results and Qualitative Study
 The experimental results demonstrate that the model trained using Gaussian noise exhibits a better anomaly detection performance to models trained with other pseudo-anomaly generator.
 To further validate the results, we present detailed histograms for the anomaly detection scores obtained across 15 individual categories as well as the overall dataset.
-These histograms were generated under the condition where the Anomaly Rate Quotient (ARQ) is set to 0.006, using RD++ as our baseline framework.
+These histograms are generated under the condition where the Anomaly Rate Quotient (ARQ) is set to 0.006, using RD++ as our baseline framework.
 Each histogram in  figs.
-3 and 4 provides a clear visualization of the distribution of detection scores, overlaid with a Gaussian fitting curve for enhanced interpretability.
+6 and 7 provides a clear visualization of the distribution of detection scores, overlaid with a Gaussian fitting curve for enhanced interpretability.
 This fitting curve highlights the distribution, allowing for a more effective comparison of the model’s performance across different categories and the overall dataset, further solidifying the reliability of our evaluation metrics.
+C.4. Future Directions with GLASS-like Pseudo-Anomaly Generators
+While Gaussian noise has proven effective as a preliminary pseudo-anomaly generator, future work could explore more advanced methods inspired by the GLASS approach (Chen et al., 2024).
+The GLASS method introduces a structured way to generate pseudo-anomalies by leveraging geometric transformations and localized pattern synthesis, providing greater control over anomaly characteristics.
+By incorporating concepts from GLASS, we envision a Gaussian Noise-Enhanced GLASS-like Generator, where Gaussian noise could act as the foundational perturbation, and additional constraints or transformations could be applied to shape the pseudo-anomalies in a contextually relevant manner.
+For instance, augmenting Gaussian noise with spatial or intensity correlations reflective of real anomalies could enhance its realism and improve model robustness in detecting diverse anomalies.
+Such an approach would align well with our COAD framework, enabling a more versatile training process that accounts for both unstructured and structured anomaly patterns.
+By systematically integrating these future developments, we aim to refine our anomaly detection strategy, bridging the gap between theoretical advancements and practical applications.
+Appendix D Additional Experimental Results on VisA Dataset
+In this section, we provide additional quantitative results for the multi-class anomaly detection task evaluated on the VisA dataset, which complements the results presented in the main paper.
+Detailed performance comparisons across different methods are summarized in Table 9.
+Appendix E Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
+To further validate the effectiveness of our methods, we provide a qualitative comparison of anomaly detection results across 15 distinct categories using baselines such as RD, RD++, UniAD, DiAD, and the proposed COAD framework in  figs.
+8, 9, 10 and 11.
+Each subfigure demonstrates: (from left to right) the original image, the ground truth anomaly mask, the result from the baseline method, and the result after applying COAD.
+The comparison shows that COAD effectively reduces both false positive rates (FPR) and false negative rates (FNR), accurately capturing anomaly regions while minimizing misclassifications.
+These results are consistent with the improvements reflected in the earlier quantitative data in  tables 4 and 5, where COAD consistently outperforms its counterparts in image-level and pixel-level AUROC metrics, highlighting COAD’s capability to outperform the baselines.
+By leveraging controllable overfitting, COAD enhances the model’s sensitivity to true anomalies while maintaining its generalization capabilities.
+The reduction in FPR minimizes the misclassification of normal regions as anomalies, while the decreased FNR highlights the model’s ability to capture subtle anomalies that baseline methods often miss.
+These qualitative and quantitative improvements collectively demonstrate the robustness and effectiveness of COAD in advancing anomaly detection tasks.
+Additionally, the sharper distinction between normal and anomalous regions demonstrates the framework’s capacity to refine boundary precision, further enhancing detection reliability.
 (a) Bottle (Anomaly)
@@
 (o) Zipper (Anomaly)
-Figure 3: Histograms for anomaly scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
+Figure 6. Histograms for anomaly scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
 (a) Bottle (Normal)
@@
 (o) Zipper (Normal)
-Figure 4: Histograms for normal scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
-8.4 Future Directions with GLASS-like Pseudo-Anomaly Generators
-While Gaussian noise has proven effective as a preliminary pseudo-anomaly generator, future work could explore more advanced methods inspired by the GLASS approach [4].
-The GLASS method introduces a structured way to generate pseudo-anomalies by leveraging geometric transformations and localized pattern synthesis, providing greater control over anomaly characteristics.
-By incorporating concepts from GLASS, we envision a Gaussian Noise-Enhanced GLASS-like Generator, where Gaussian noise could act as the foundational perturbation, and additional constraints or transformations could be applied to shape the pseudo-anomalies in a contextually relevant manner.
-For instance, augmenting Gaussian noise with spatial or intensity correlations reflective of real anomalies could enhance its realism and improve model robustness in detecting diverse anomalies.
-Such an approach would align well with our COAD framework, enabling a more versatile training process that accounts for both unstructured and structured anomaly patterns.
-By systematically integrating these future developments, we aim to refine our anomaly detection strategy, bridging the gap between theoretical advancements and practical applications.
-9 Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
-To further validate the effectiveness of our methods, we provide a qualitative comparison of anomaly detection results across 15 distinct categories using baselines such as RD, RD++, UniAD, DiAD, and the proposed COAD framework in  Figs.
-5, 6, 7 and 8.
-Each subfigure demonstrates: (from left to right) the original image, the ground truth anomaly mask, the result from the baseline method, and the result after applying COAD.
-The comparison shows that COAD effectively reduces both false positive rates (FPR) and false negative rates (FNR), accurately capturing anomaly regions while minimizing misclassifications.
-These results are consistent with the improvements reflected in the earlier quantitative data in  Tabs.
-3 and 4, where COAD consistently outperforms its counterparts in image-level and pixel-level AUROC metrics, highlighting COAD’s capability to outperform the baselines.
-By leveraging controllable overfitting, COAD enhances the model’s sensitivity to true anomalies while maintaining its generalization capabilities.
-The reduction in FPR minimizes the misclassification of normal regions as anomalies, while the decreased FNR highlights the model’s ability to capture subtle anomalies that baseline methods often miss.
-These qualitative and quantitative improvements collectively demonstrate the robustness and effectiveness of COAD in advancing anomaly detection tasks.
-Additionally, the sharper distinction between normal and anomalous regions demonstrates the framework’s capacity to refine boundary precision, further enhancing detection reliability.
+Figure 7. Histograms for normal scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
 (a) Bottle
@@
 (o) Zipper
-Figure 5: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using the Baseline Method RD and COAD.
+Figure 8.
+Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using the Baseline Method RD and COAD.
 Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
 (o) Zipper
-Figure 6: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method RD++ and COAD.
+Figure 9.
+Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method RD++ and COAD.
 Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
 (o) Zipper
-Figure 7: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method UniAD and COAD.
+Figure 10.
+Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method UniAD and COAD.
 Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
 (o) Zipper
-Figure 8: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method DiAD and COAD.
+Figure 11.
+Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method DiAD and COAD.
 Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
-10 Future Work
-In future work, we plan to refine our pseudo-anomaly generators by incorporating advanced methods in  Sec. 8.4, enabling the generation of more structured and domain-relevant pseudo-anomalies.
+Appendix F Future Work
+In future work, we plan to refine our pseudo-anomaly generators by incorporating advanced methods in  section C.4, enabling the generation of more structured and domain-relevant pseudo-anomalies.
 Moreover, we intend to explore the application of the COAD framework in other domains, particularly in areas such as medical imaging, autonomous driving, and forgery detection.